### v1

# Evidence-based multimodal fusion on structured EHRs and free-text notes for ICU outcome prediction

ABSTRACT
Objective: Accurate Intensive Care Unit (ICU) outcome prediction is critical for improving patient treatment quality and ICU resource allocation. Existing research mainly focuses on structured data and lacks effective frameworks to integrate clinical notes from heterogeneous electronic health records (EHRs). This study aims to explore a multimodal framework based on evidence theory that can effectively combine heterogeneous structured EHRs and free-text notes for accurate and reliable ICU outcome prediction.
Materials and Methods: We proposed an evidence-based multimodal fusion framework to predict ICU outcomes, including mortality and prolonged length of stay (PLOS), by utilizing both structured EHR data and free-text notes from the MIMIC-III database. We compare the performance against baseline models that use only structured EHRs, free-text notes, or existing multimodal approaches.
Results: The results demonstrate that the evidence-based multimodal fusion model achieved both accurate and reliable prediction. Specifically, it outperformed the best baseline by 1.05%/1.02% in BACC, 9.74%/6.04% in F1 score, 1.28%/0.9% in AUROC, and 6.21%/2.68% in AUPRC for predicting mortality and PLOS, respectively. Additionally, it improved the reliability of the predictions with a 26.8%/15.1% reduction in the Brier score and a 25.0%/13.3% reduction in negative log-likelihood.
Conclusion: This study demonstrates that the evidence-based multimodal fusion framework can serve as a strong baseline for predictions using structured EHRs and free-text notes. It effectively reduces false positives, which can help improve the allocation of medical resources in the ICU. This framework can be further applied to analyze multimodal EHRs for other clinical tasks.
Keywords: Multimodal learning, Evidence fusion, Dempster-Shafer theory, ICU outcome prediction, Electronic health records
INTRODUCTION
The Intensive Care Unit (ICU) is a specialized hospital ward that offers comprehensive and continuous care to critically ill patients. As the population of critically ill patients grows, the demand for ICUs has risen significantly, placing strain on already limited and costly intensive care resources 1, especially during public health crises like the COVID-19 pandemic, when hospitals face an overwhelming surge of patients.
Due to the limited availability of intensive care resources, researchers have emphasized the necessity of predicting ICU outcomes such as mortality rates and prolonged lengths of stay (PLOS). Accurate predictions can help in the efficient allocation of medical resources for patients in need and reduce unnecessary expenses without compromising patient care. Furthermore, they are crucial for healthcare providers in making informed decisions about patient care strategies and providing early interventions to patients at high risk of adverse outcomes 2, 3.
Over the past two decades, the adoption of electronic ICU technology has enabled the collection of extensive data on ICU patients, creating new opportunities for developing advanced methods to predict ICU outcomes. Most previous research has concentrated on modeling ICU outcome predictions using structured EHR data 4, 5, 6, which often captures only a portion of clinical information. It may miss out on the rich contextual information that unstructured EHR data (such as nursing notes, patient narratives, and imaging reports) can provide. Natural language processing (NLP) techniques have been well explored to extract valuable insights from unstructured free-text EHR data 7, 8, 9. Therefore, effective multimodal learning algorithms are essential to integrate heterogeneous EHRs for better ICU outcome prediction. Recently, deep learning-based multimodal models have been proven to combine structured EHR data and free-text data at the deep feature level for clinical outcome predictions 10, 11, 12. While those methods have increased the prediction accuracy, the question regarding the reliability of those models in noisy and unstable clinical situations still exists.
In this study, we hypothesize that an effective multimodal deep neural network that incorporates both structured EHR data and free-text notes can improve the prediction of ICU outcomes. Our objective is to propose and evaluate a novel evidence-based multimodal framework for ICU outcome prediction using multimodal EHR data. Unlike traditional methods that rely on data or feature embeddings, our framework integrates information based on evidence derived from different modalities. Specifically, we use state-of-the-art deep neural networks for single-modality feature extraction: ResNet/Transformer-based models for structured EHR data and pre-trained language models for free-text EHR data. The extracted features are independently mapped into evidence and then combined in the evidence space. Experimental results on the MIMIC-III database for mortality and prolonged length of stay (PLOS) predictions demonstrate the effectiveness of the proposed model. The code implementation is available on GitHub repository (https://github.com/yuchengruan/evid_multimodal_ehr)
MATERIALS AND METHODS
Overview
In this section, we demonstrate the cohort, patient features, prediction outcomes, and proposed evidence-based multimodal fusion framework in combining structured EHRs and free-text notes for ICU outcome prediction.
Cohort
MIMIC-III (Medical Information Mart for Intensive Care III) is a large, publicly available database containing de-identified health records from patients in critical care units at Beth Israel Deaconess Medical Center (from the US) between 2001 and 2012 13. We collected structured EHR data and free-text clinical notes from the database. Patients were excluded if they (1) were under 18 years of age at admission and (2) had incomplete length of stay or mortality data. For patients with multiple ICU stays, we considered only the first.
Patient features
Patient features include both structured EHR data and unstructured free-text notes. In this section, we demonstrate the patient features in our study and provide details about the data preprocessing steps.
Structured EHR data
The structured data were collected during patients’ ICU stays and included demographic information, vital signs and laboratory tests, medical treatments, and comorbidities. For demographic information, the patient’s age, gender, weight, ethnicity, and admission type at the time of admission were included in the study. Vital signs/lab tests are the most crucial health indicators, easily measured using non-invasive equipment, and are readily understood by all healthcare professionals. For each variable considered, we used the first value recorded within 24 hours of admission time. We then excluded any variables with $\geq 50\%$ missingness rate; this resulted in the inclusion of only heart rate among vital signs features alongside 19 lab test features, such as blood urea nitrogen, eosinophil count, and lymphocyte count, over the same period. All vital sign/lab test features were numerical variables. Medical treatments, which include services and interventions provided to patients and recorded in digital systems, were also analyzed. Treatments such as sedatives, statins, diuretics, antibiotics, ventilation, and vasopressors were included, with each treatment feature coded as a binary variable, indicating whether the patient received the treatment. Comorbidities refer to the presence of additional medical conditions, which play a role in decision-making models. In this study, comorbidities such as hypertension, diabetes, alcohol abuse, cerebrovascular accident (stroke), congestive heart failure, and ischemic heart disease were included, all represented as binary variables.
All categorical features were encoded using one-hot encoding, and numerical features were normalized. Missing data were addressed by imputing the mean for continuous features and the mode for categorical features, ensuring data consistency. Eventually, the structured data contained 41 features.
Free-text EHR notes
Free-text notes contain a rich repository of clinical information about observations, assessments, and the overall clinical picture, which structured data often fails to capture. Furthermore, they provide an important context for interpreting structured data. For instance, while lab results may indicate abnormal values, free-text notes can clarify the relevance of these results by considering the patient’s history, comorbidities, or specific circumstances at the time of testing. As a result, NLP techniques, especially pre-trained LLM, can be applied to these notes to gain deeper insights for data-driven predictions. In this study, we focus on Nursing, Nursing/Other, Physician, and Radiology notes, as these comprise the majority of clinical documentation and are frequently recorded in the MIMIC-III database 10. We extracted only the first 24 hours of notes for each admission to facilitate early outcome prediction.
[…]
Prediction Outcomes
In this study, we focus on two ICU prediction outcomes: mortality and prolonged length of stay.
Mortality
Mortality is widely acknowledged as a critical outcome in ICUs. The primary objective of this task is to determine whether a patient is likely to die during their hospital stay. Accurate predictions enable the early identification of high-risk patients and support the efficient allocation of ICU resources. This prediction task is typically framed as a binary classification problem, with the label indicating the occurrence or absence of a death event.
Prolonged length of stay
Length of stay refers to the duration between a patient’s admission to and discharge from the ICU. In this study, we aim to predict prolonged length of stay (PLOS), defined as a stay exceeding 7 days 14, 15, 10. Prolonged ICU stays are often linked to severe illnesses, complications, and increased mortality. Moreover, they place considerable pressure on hospital resources by reducing the availability of ICU beds and specialized personnel. Efficient management of ICU LOS not only improves patient outcomes but also enhances the overall effectiveness of healthcare systems. This problem is framed as a binary classification task.
Model architecture
The overview of the proposed framework is illustrated in Figure 1. In general, we propose to generate the modality-level evidence for both structured data and free-text notes using the evidence mapping module and fuse the modality-level evidence for final prediction.
Figure 1: The overview of our proposed framework. EM: evidence mapping and EF: evidence fusion
Evidence mapping (EM)
Denœux 16 proposed an evidential neural network (ENN) that maps imperfect (uncertain, imprecise, or noise) input features into degrees of belief and ignorance (uncertainty) under the framework of Dempster-Shafer (DS) theory 17, 18. The essential concept of ENN is to consider each prototype as a piece of evidence, which is discounted based on its distance from the input vector. The evidence from different prototypes is then aggregated using Dempster’s combination rule. Inspired by ENN and its promising adoptions in medical data analysis 19, 20, 21, 22, we propose incorporating ENN as an evidence mapping module with the state-of-the-art encoders to generate evidence for structured EHRs and free-text notes. Further details about DS theory and evidence mapping can be found in Appendix A and B. Given modality level input, the evidence mapping module can output the evidence for each class as well as the uncertainty regarding this prediction.
Structured data evidence mapping
To produce modality evidence for structured data, we initially used a structured data encoder to extract deep features (the output dimension is set to 32). We considered three popular encoders to extract the embeddings: MLP, ResNet, FT-Transformer 23 (see Baselines section for more details). Subsequently, we introduced an evidence mapping module to transform the deep features into evidence embeddings for structured data.
Free-text notes evidence mapping
Similarly, we utilized pre-trained language models to extract the deep features from clinical notes, on which we developed an evidence mapping module to produce the evidence of the modality. Our primary analysis focused on pre-trained architectures similar to BERT. Accordingly, three BERT-based architectures were evaluated: BERT 24, BioBERT 24, Clinical BERT 25 (see Baselines section for more details). To minimize computational overhead while maintaining high predictive performance, we froze the pre-trained language models and fine-tuned an additional layer (128 hidden units) on top in model training.
Evidence fusion (EF)
Based on the modality-level evidence obtained from structured EHR and free-text notes, we developed the evidence fusion module based on Dempster’s combination rule to generate the final evidence for decision-making. Further details about the fusion of EHR and free-text notes evidence can be found in Appendix C.
Augmented learning algorithm
We optimize the proposed framework using an augmented learning algorithm, which includes two types of optimization objectives: (1) main objective and (2) auxiliary objective. The main objective is to optimize predictive performance based on transformed evidence as the primary loss function. Additionally, two auxiliary cross-entropy losses are incorporated to enhance the feature representation capability of the independent encoders for the two modalities, as the evidence mapping module performs more effectively with high-quality representations. More details can be found in the Appendix D.
Evaluation and experiment
Baselines
To comprehensively evaluate the effectiveness of our proposed fusion framework, we employed three types of classification models as baselines: (1) models with structured data only, (2) models with free-text notes only, and (3) existing multimodal models based on both data. The configurations for the structured data models are as follows:
Random Forest 26: A decision tree-based ensemble learning method for making predictions.
MLP 23: A fundamental neural network with fully connected layers to encode structured data and serves as a reliable baseline. The model was configured with 3 layers, 32 hidden units, and a dropout rate of 0.1.
ResNet 23: Because of the success of ResNet in computer vision 27, it has also been adapted for structured data modeling. Specifically, the main building block is simplified by providing a direct path from input to output. The configuration in our study has 3 residual blocks, 32 hidden units, and a dropout rate of 0.1.
FT-Transformer 23: It converts all categorical and numerical features into embeddings, which are then processed through a couple of Transformer layers. It has demonstrated superior performance as a structured data encoder across various tasks. The model configuration has 3 Transformer layers with 192 hidden units, 8 attention heads, and a dropout rate of 0.2.
As our focus is on pre-trained language models similar to BERT, we employed three BERT-based text classification models in this study:
BERT 24: a pre-trained large language model English corpus using a self-supervised approach. Through masked language modeling and next-sentence prediction objectives, the model learns internal representations of the English language, enabling it to extract features useful for downstream tasks. In this study, we utilized the google-bert/bert-base-uncased model from Huggingface 28 for feature extraction.
BioBERT 29: a variant of BERT pre-trained on biomedical literature, such as PubMed abstracts, and is optimized to perform more effectively on biomedical NLP tasks. In this study, we used dmis-lab/biobert-v1.1 from Huggingface as the extraction model.
Clinical BERT 25: A fine-tuned version of BERT on clinical notes from the MIMIC-III database, making it well-suited for handling medical terminology and clinical narratives to enhance performance on clinical tasks. In our study, we used emilyalsentzer/Bio_ClinicalBERT from the Huggingface transformer library for extracting embeddings from clinical notes.
We also implemented a Multimodal approach for the ICU outcome prediction from 12: it concatenated the structured EHRs with the extracted embeddings from free-text notes, which were then passed into 2 fully connected layers for prediction. It supports different text encoders, and in this study, we used BERT, BioBERT, and Clinical BERT for fair comparisons
Implementation details
The dataset was randomly divided with 60% for training, 20% for validation, and 20% for testing. For model training, a mini-batch size of 32 was used, and the maximum number of epochs was set to 150, with early stopping applied. Each model was trained five times with different random seeds, and we reported the average results along with the standard error of the metrics to ensure statistical reliability. Hyperparameters were optimized for all baseline models to achieve the best results.
All compared models were implemented using Scikit-learn 30, PyTorch 31, and Hugging Face’s Transformers library 28 in Python 3.8.19. The MLP, ResNet, and FT-Transformer models were built using the original source code on GitHub (https://github.com/yandex-research/rtdl-revisiting-models). The Multimodal approach was modified based on the code on Github (https://github.com/WeiChunLin/Bio_Clinical_BERT_Multimodal_Model).
Model evaluation
For comprehensive model evaluation and comparison, we reported the following metrics: precision, recall, specificity, negative predictive value (NPV), balanced accuracy (BACC), F1 score, area under the receiver operating characteristic curve (AUROC), area under the precision-recall curve (AUPRC), Brier score, and negative log-likelihood (NLL). Precision, recall, specificity, and NPV assess model performance for positive or negative samples. In contrast, BACC, F1 score, AUROC, and AUPRC provide a more holistic evaluation of predictive performance. Meanwhile, the Brier score and NLL measure the prediction reliability.
Other fusion settings
In addition to fusing evidence from modalities (structured EHRs and free-text notes), we evaluated our framework under two additional fusion settings: data types and data sources. The illustration is shown in Figure 2.
Figure 2: The illustration of different fusion settings.
Data types: Features can be grouped based on their data types. In our study, we further divided the structured data into two types: numerical and categorical data.
Data Sources: Features can also be grouped according to their sources. In our case, the structured data was split into four sources: demographics, vital signs/lab tests, medical treatments, and comorbidities. Similarly, we categorized the clinical notes into four types: Nursing, Nursing/Other, Physician, and Radiology notes. For each group, we generated evidence using individual encoders and evidential layers. These pieces of evidence were then aggregated through the EF module to produce the final evidence for decision-making.
RESULTS
Data description
Table 1 shows the descriptive characteristics of the patients in the study cohort. Our cohort includes 38469 patients in total under the inclusion criteria, in which 4540 (11.8%) patients were identified as dead during the stay while 5220 (13.6%) patients were identified as having prolonged length of stay. The patient demographic showed that the majority of patients were white, and male patients were slightly more than the female. Most of the admitted patients in the ICU were identified as emergency. Over half (51.2%) of patients received mechanical ventilation during their ICU stay and hypertension (38.8%) and congestive heart failure (31.2%) were among the most common patient comorbidities.
Patient chracteristic Mortality PLOS
Yes (N=4540) No (N=33928) Yes (N=5220) No (N=33248)
Age 68.6 (15.1) 61.6 (16.9) 62.9 (16.3) 62.3 (16.9)
Gender
  Male 2399 (52.8%) 19378 (57.1%) 2943 (56.4%) 18834 (56.6%)
  Female 2141 (47.2%) 14550 (42.9%) 2277 (43.6%) 14414 (43.4%)
Weight 79.1 (22.7) 83.0 (23.1) 84.6 (24.3) 82.2 (22.9)
Ethnicity
  White 3100 (68.3%) 24343 (71.7%) 3661 (70.1%) 23782 (71.5%)
  Black 260 (5.7%) 2688 (7.9%) 354 (6.8%) 2594 (7.8%)
  Asian 106 (2.3%) 1166 (3.4%) 153 (2.9%) 1101 (3.3%)
  Hispanic 88 (1.9%) 803 (2.4%) 104 (2.0%) 805 (2.4%)
  Other/Unknown 986 (21.7%) 4928 (14.5%) 948 (18.2%) 4966 (14.9%)
Admission type
  Emergency 4240 (93.4%) 27062 (79.8%) 4480 (85.8%) 26822 (80.7%)
  Elective 165 (3.6%) 5911 (17.4%) 533 (10.2%) 5543 (16.7%)
  Urgent 135 (3.0%) 955 (2.8%) 207 (4.0%) 883 (2.7%)
Heart rate 91.4 (22.0) 86.8 (18.7) 91.5 (21.0) 86.7 (18.8)
APTT 39.7 (27.1) 34.7 (21.6) 37.2 (23.8) 35.0 (22.2)
BUN 34.1 (25.1) 23.8 (19.0) 28.1 (22.3) 24.5 (19.7)
Eosinophil 1.3 (3.1) 1.5 (1.9) 1.2 (1.7) 1.5 (2.1)
Lymphocytes 12.1 (11.9) 15.5 (11.4) 12.5 (10.8) 15.4 (11.6)
Neutrophils 77.3 (17.7) 76.6 (13.9) 77.8 (15.2) 76.5 (14.4)
RDW 15.5 (2.4) 14.5 (1.9) 14.9 (2.1) 14.6 (2.0)
Bicarbonate 22.6 (5.7) 24.3 (4.4) 23.6 (5.2) 24.1 (4.5)
Chloride 103.3 (7.3) 104.1 (6.0) 104.0 (6.7) 104.0 (6.1)
Creatinine 1.6 (1.4) 1.3 (1.4) 1.4 (1.5) 1.3 (1.4)
Hemoglobin 11.3 (2.3) 11.8 (2.3) 11.6 (2.2) 11.7 (2.3)
Mean cell volume 91.5 (7.7) 89.4 (6.6) 90.3 (7.1) 89.6 (6.7)
Platelet count 227.2 (132.2) 239.0 (112.3) 231.8 (119.8) 238.5 (114.0)
Potassium 4.3 (0.9) 4.2 (0.7) 4.2 (0.8) 4.2 (0.7)
Sodium 138.2 (6.2) 138.6 (4.6) 138.7 (5.2) 138.5 (4.7)
PT 17.4 (10.6) 15.1 (6.8) 16.0 (8.2) 15.3 (7.3)
INR 1.8 (2.2) 1.4 (1.3) 1.6 (1.9) 1.4 (1.4)
WBC 14.3 (16.5) 11.5 (8.8) 12.9 (8.8) 11.6 (10.2)
PLR 39.9 (56.7) 30.4 (47.4) 38.4 (57.5) 30.6 (47.3)
NLR 13.9 (15.6) 9.9 (13.7) 13.0 (15.2) 10.1 (13.8)
Sedatives 1263 (27.8%) 9114 (26.9%) 2360 (45.2%) 8017 (24.1%)
Statin 405 (8.9%) 5164 (15.2%) 556 (10.7%) 5013 (15.1%)
Diuretic 567 (12.5%) 5435 (16.0%) 865 (16.6%) 5137 (15.5%)
Antibiotics 952 (21.0%) 4836 (14.3%) 1122 (21.5%) 4666 (14.0%)
Ventilation 2326 (51.2%) 10923 (32.2%) 3105 (59.5%) 10144 (30.5%)
Vasopressor 1538 (33.9%) 6340 (18.7%) 1639 (31.4%) 6239 (18.8%)
Hypertension 1760 (38.8%) 16221 (47.8%) 2199 (42.1%) 15782 (47.5%)
Diabetes 1107 (24.4%) 9249 (27.3%) 1401 (26.8%) 8955 (26.9%)
Alcohol abuse 202 (4.4%) 1554 (4.6%) 333 (6.4%) 1423 (4.3%)
CVA 309 (6.8%) 1139 (3.4%) 364 (7.0%) 1084 (3.3%)
CHF 1416 (31.2%) 8711 (25.7%) 1867 (35.8%) 8260 (24.8%)
IHD 1276 (28.1%) 12390 (36.5%) 1634 (31.3%) 12032 (36.2%)
Table 1: Characteristics of structured features in the patient cohort. For categorical features, the number of instances in each category is reported along with the percentage. For continuous features, the mean and standard deviation are reported in the study.
Model performance
Table 2 presents the comparison of model performance under predictive evaluation metrics (BACC, F1 score, AUROC, and AUPRC) and reliability evaluation metrics (Brier score and NLL) for both mortality prediction (2(a)) and PLOS (2(b)) prediction tasks. The comparison results under precision, recall, specificity, and NPV can be found in Appendix E.
[…]
Random Forest x 0.7172$\pm$0.0019 0.3820$\pm$0.0018 0.7988$\pm$0.0010 0.3766$\pm$0.0030 0.1891$\pm$0.0006 0.5639$\pm$0.0015
MLP x 0.7486$\pm$0.0003 0.4043$\pm$0.0019 0.8326$\pm$0.0011 0.4429$\pm$0.0033 0.1677$\pm$0.0015 0.4937$\pm$0.0037
ResNet x 0.7521$\pm$0.0011 0.4113$\pm$0.0007 0.8350$\pm$0.0006 0.4468$\pm$0.0016 0.1654$\pm$0.0006 0.4895$\pm$0.0015
FT-Transformer x 0.7592$\pm$0.0025 0.4166$\pm$0.0021 0.8426$\pm$0.0013 0.4577$\pm$0.0029 0.1648$\pm$0.0014 0.4832$\pm$0.0037
BERT x 0.6300$\pm$0.0014 0.2814$\pm$0.0019 0.6777$\pm$0.0014 0.2037$\pm$0.0012 0.2302$\pm$0.0070 0.6523$\pm$0.0159
BioBERT x 0.6164$\pm$0.0043 0.2759$\pm$0.0021 0.6695$\pm$0.0016 0.2181$\pm$0.0009 0.2214$\pm$0.0021 0.6326$\pm$0.0048
Clinical Bert x 0.6614$\pm$0.0007 0.3080$\pm$0.0009 0.7240$\pm$0.0001 0.2928$\pm$0.0008 0.2132$\pm$0.0030 0.6086$\pm$0.0070
Multimodal (BERT) x x 0.7531$\pm$0.0016 0.4079$\pm$0.0031 0.8398$\pm$0.0006 0.4596$\pm$0.0021 0.1691$\pm$0.0030 0.4992$\pm$0.0077
Multimodal (BioBERT) x x 0.7558$\pm$0.0007 0.4073$\pm$0.0027 0.8362$\pm$0.0006 0.4570$\pm$0.0032 0.1725$\pm$0.0028 0.5084$\pm$0.0088
Multimodal (Clinical Bert) x x 0.7584$\pm$0.0005 0.4218$\pm$0.0019 0.8404$\pm$0.0004 0.4686$\pm$0.0026 0.1606$\pm$0.0022 0.4792$\pm$0.0046
Ours (MLP-BERT) x x 0.7525$\pm$0.0030 0.4466$\pm$0.0042 0.8406$\pm$0.0010 0.4764$\pm$0.0020 0.1345$\pm$0.0050 0.4041$\pm$0.0128
Ours (MLP-BioBERT) x x 0.7492$\pm$0.0042 0.4493$\pm$0.0060 0.8408$\pm$0.0012 0.4758$\pm$0.0009 0.1301$\pm$0.0062 0.3927$\pm$0.0159
Ours (MLP-Clinical Bert) x x 0.7467$\pm$0.0022 0.4629$\pm$0.0033 0.8465$\pm$0.0008 0.4935$\pm$0.0011 0.1176$\pm$0.0027 0.3594$\pm$0.0066
Ours (ResNet-BERT) x x 0.7581$\pm$0.0037 0.4404$\pm$0.0038 0.8415$\pm$0.0011 0.4688$\pm$0.0014 0.1435$\pm$0.0052 0.4296$\pm$0.0133
Ours (ResNet-BioBERT) x x 0.7546$\pm$0.0039 0.4438$\pm$0.0053 0.8424$\pm$0.0012 0.4671$\pm$0.0014 0.1392$\pm$0.0061 0.4187$\pm$0.0157
Ours (ResNet-Clinical Bert) x x 0.7580$\pm$0.0023 0.4472$\pm$0.0042 0.8474$\pm$0.0006 0.4899$\pm$0.0015 0.1387$\pm$0.0043 0.4178$\pm$0.0110
Ours (FT-Trans-BERT) x x 0.7634$\pm$0.0012 0.4432$\pm$0.0060 0.8485$\pm$0.0007 0.4797$\pm$0.0029 0.1450$\pm$0.0049 0.4345$\pm$0.0129
Ours (FT-Trans-BioBERT) x x 0.7610$\pm$0.0015 0.4507$\pm$0.0007 0.8486$\pm$0.0011 0.4796$\pm$0.0027 0.1358$\pm$0.0014 0.4095$\pm$0.0045
Ours (FT-Trans-Clinical Bert) x x 0.7672$\pm$0.0032 0.4541$\pm$0.0032 0.8534$\pm$0.0012 0.4977$\pm$0.0011 0.1390$\pm$0.0043 0.4192$\pm$0.0112
(a) Mortality prediction
[…]
Random Forest x 0.6680$\pm$0.0025 0.3492$\pm$0.0038 0.7270$\pm$0.0018 0.2980$\pm$0.0029 0.2350$\pm$0.0009 0.6649$\pm$0.0018
MLP x 0.6845$\pm$0.0009 0.3736$\pm$0.0011 0.7528$\pm$0.0010 0.3353$\pm$0.0010 0.1950$\pm$0.0018 0.5739$\pm$0.0044
ResNet x 0.6851$\pm$0.0017 0.3774$\pm$0.0018 0.7532$\pm$0.0004 0.3366$\pm$0.0021 0.1929$\pm$0.0012 0.5702$\pm$0.0027
FT-Transformer x 0.6956$\pm$0.0011 0.3790$\pm$0.0021 0.7674$\pm$0.0006 0.3404$\pm$0.0021 0.1947$\pm$0.0031 0.5704$\pm$0.0079
BERT x 0.6087$\pm$0.0011 0.2960$\pm$0.0016 0.6507$\pm$0.0015 0.2262$\pm$0.0009 0.2316$\pm$0.0043 0.6557$\pm$0.0091
BioBERT x 0.6057$\pm$0.0016 0.2940$\pm$0.0012 0.6487$\pm$0.0014 0.2214$\pm$0.0012 0.2297$\pm$0.0029 0.6518$\pm$0.0063
Clinical Bert x 0.6420$\pm$0.0016 0.3261$\pm$0.0015 0.6946$\pm$0.0009 0.2661$\pm$0.0006 0.2225$\pm$0.0031 0.6350$\pm$0.0069
Multimodal (BERT) x x 0.6903$\pm$0.0018 0.3739$\pm$0.0012 0.7625$\pm$0.0006 0.3427$\pm$0.0007 0.2017$\pm$0.0028 0.5960$\pm$0.0082
Multimodal (BioBERT) x x 0.6905$\pm$0.0022 0.3731$\pm$0.0025 0.7606$\pm$0.0007 0.3364$\pm$0.0009 0.1974$\pm$0.0026 0.5888$\pm$0.0043
Multimodal (Clinical Bert) x x 0.6933$\pm$0.0014 0.3714$\pm$0.0023 0.7648$\pm$0.0009 0.3544$\pm$0.0018 0.2027$\pm$0.0056 0.5871$\pm$0.0136
Ours (MLP-BERT) x x 0.6878$\pm$0.0016 0.3868$\pm$0.0030 0.7580$\pm$0.0004 0.3499$\pm$0.0015 0.1773$\pm$0.0022 0.5288$\pm$0.0063
Ours (MLP-BioBERT) x x 0.6866$\pm$0.0022 0.3905$\pm$0.0038 0.7573$\pm$0.0010 0.3461$\pm$0.0023 0.1720$\pm$0.0031 0.5172$\pm$0.0076
Ours (MLP-Clinical Bert) x x 0.6870$\pm$0.0012 0.3952$\pm$0.0032 0.7633$\pm$0.0008 0.3557$\pm$0.0021 0.1637$\pm$0.0054 0.4943$\pm$0.0145
Ours (ResNet-BERT) x x 0.6931$\pm$0.0013 0.3887$\pm$0.0040 0.7655$\pm$0.0011 0.3584$\pm$0.0007 0.1815$\pm$0.0045 0.5398$\pm$0.0107
Ours (ResNet-BioBERT) x x 0.6896$\pm$0.0011 0.3918$\pm$0.0037 0.7647$\pm$0.0009 0.3532$\pm$0.0010 0.1752$\pm$0.0048 0.5241$\pm$0.0123
Ours (ResNet-Clinical Bert) x x 0.6974$\pm$0.0011 0.4019$\pm$0.0015 0.7705$\pm$0.0009 0.3639$\pm$0.0010 0.1694$\pm$0.0018 0.5082$\pm$0.0042
Ours (FT-Trans-BERT) x x 0.7005$\pm$0.0010 0.3809$\pm$0.0024 0.7725$\pm$0.0006 0.3524$\pm$0.0011 0.1927$\pm$0.0039 0.5617$\pm$0.0100
Ours (FT-Trans-BioBERT) x x 0.6986$\pm$0.0011 0.3847$\pm$0.0030 0.7701$\pm$0.0012 0.3504$\pm$0.0014 0.1832$\pm$0.0045 0.5376$\pm$0.0113
Ours (FT-Trans-Clinical Bert) x x 0.7027$\pm$0.0011 0.3942$\pm$0.0020 0.7743$\pm$0.0006 0.3575$\pm$0.0014 0.1760$\pm$0.0018 0.5205$\pm$0.0038
(b) PLOS prediction
Table 2: The model comparison for mortality (a) and PLOS prediction (b) tasks.
Overall, our framework demonstrated superior performance across both tasks. In the mortality prediction task, our framework using MLP and Clinical BERT as the backbone achieved the highest F1 score (0.4629) and the lowest Brier score (0.1176) and NLL (0.3594). Furthermore, with FT-Transformer and Clinical BERT, it achieved the highest balanced accuracy (BACC) of 0.7672, AUROC of 0.8534, and AUPRC of 0.4977. Compared to the best baseline models, our framework improved predictive performance by approximately 1.05% in BACC, 9.74% in F1 score, 1.28% in AUROC, and 6.21% in AUPRC. Additionally, it significantly enhanced prediction reliability, with reductions of about 26.8% in Brier score and 25.0% in NLL.
In the PLOS prediction task, our framework demonstrated exceptional performance across all key predictive evaluation metrics, achieving a BACC of 0.7027, an F1 score of 0.4019, an AUROC of 0.7743, and an AUPRC of 0.3639. Compared to the best baseline model, our framework showed notable improvements: a 1.02% increase in BACC, a 6.04% increase in F1 score, a 0.9% increase in AUROC, and a 2.68% increase in AUPRC. Additionally, it enhanced prediction reliability with a 15.1% reduction in Brier score and a 13.3% reduction in NLL.
Models that relied solely on structured data or free-text notes underperformed compared to our proposed framework in both predictive and reliability evaluation metrics. While the multimodal approach showed marginally better performance than other baselines in the mortality prediction task, it delivered weaker results in the PLOS prediction task.
Evaluation on different fusion settings
Figure 3 illustrates the evaluation of our framework across three fusion settings applied to different backbone models for mortality prediction. For the corresponding results in the PLOS prediction task, please refer to Appendix F. Both tasks exhibited similar patterns in the analysis.
Regarding predictive evaluation metrics, the models achieved comparable BACC across the three fusion settings for both tasks. Notably, the model based on modalities outperformed those based on data types and data sources in F1, AUROC, and AUPRC metrics. In terms of reliability evaluation metrics, the model based on modalities demonstrated superior reliability across different backbone models.
Figure 3: The evaluation of our framework on different fusion settings for mortality prediction: (1) modalities, (2) data types, (3) data sources.
DISCUSSION
Principle findings
In this section, we present our findings across four key areas: the advantages of using free-text notes for mortality and PLOS prediction, the effectiveness of evidence-based multimodal framework, the impact of various fusion settings, and the influence of encoder selection.
Benefits of free-text notes
From Table 2, it is evident that integrating free-text notes with structured data enhances ICU outcome prediction, boosting predictive performance and reliability. Free-text notes provide information not included in structured EHR data, such as nursing details, physician documentation, and radiology reports after ICU admission. This suggests that free-text EHR notes and structured inputs can complement each other in predictive modeling, leading to improved performance.
Effectiveness of evidence-based multimodal framework
Our evidence-based framework demonstrated improved predictive performance and higher reliability compared to existing multimodal approaches. This improvement is likely due to the heterogeneous nature of EHRs, which often contain imprecise information from diverse sources. By effectively extracting and integrating informative evidence through modalities, our framework addresses these challenges.
Additionally, the current multimodal approach achieved higher recall in both tasks (refer to Appendix E). High recall reflects the model’s ability to capture the most true positive instances, often at the cost of accepting more false positives. In ICU settings, this could result in the unnecessary utilization of medical resources and equipment. In contrast, our framework strikes a better balance by demonstrating higher precision and specificity, reducing false positives. This capability is critical to appropriately allocate medical resources to ICU patients in real need.
It is also noteworthy that the improvement in BACC (1.05% for mortality prediction and 1.02% for PLOS prediction) is smaller than the improvement in F1 score (9.74% for mortality prediction and 6.04% for PLOS prediction). The F1 score, which emphasizes the balance between precision and recall, is particularly valuable for assessing improvements in identifying positive samples within imbalanced datasets, especially when positive outcomes are rare. Given that ICU outcomes are a minority in our dataset, this suggests that our framework is particularly effective at correctly identifying these critical outcomes.
Impacts of different fusion settings
To explore the performance of our framework across different fusion settings, it is evident that models based on modalities achieved higher F1 scores but lower BACCs compared to the other two fusion settings. This discrepancy arises from the metrics’ focus: the F1 score emphasizes performance on the positive class by balancing precision and recall, while balanced accuracy (BACC) provides an overall assessment of recall across all classes. Thus, models based on modalities are more effective at identifying ICU outcomes, likely due to the enhanced integration of information from independent sources facilitated by evidence theory.
This observation aligns with our analysis of data independence. Figures 4 and 5 visualize the independence of structured features and four types of free-text notes, respectively. In Figure 4, correlation coefficients confirm that features within structured data are not independent. Meanwhile, Figure 5 shows that Radiology notes form a distinct cluster, while the other types of notes overlap, indicating a lack of independence among them. Additionally, our analysis highlights a general relationship between predictive performance and reliability.
Figure 4: The pair plots and correlation coefficients of some structured features from different sources.
Figure 5: The t-SNE visualization on extracted embeddings of four different types of free-texts in EHRs: Nursing, Nursing/other, Physician, Radiology.
Influence of encoders selection
The evaluations presented in Tables 2 and 3 reveal that the choice of encoders plays a critical role in determining predictive performance. Among the models assessed, the FT-Transformer stands out as highly effective for tabular data, aligning with the findings of 23. This effectiveness can be attributed to the transformer’s capability to capture complex relationships among transformed numerical and categorical features, which enhances its predictive power.
For pre-trained language encoders, Clinical BERT demonstrates the best performance in extracting clinical information from free-text notes. Its superiority stems from its unique pre-training on clinical text from the MIMIC-III database. This specialized pre-training enables Clinical BERT to generate more informative embeddings by leveraging its pre-learned clinical knowledge and domain-specific term embeddings, resulting in improved model performance.
CONCLUSION
In this paper, we address the challenge of predicting ICU outcomes by introducing an evidence-based multimodal framework that models both structured EHR data and free-text EHR notes. Our framework transforms deep features extracted from these two modalities into evidence through evidence mapping module, which is then fused using Dempster’s rule to make final predictions. Through experiments on the MIMIC-III dataset, we demonstrate the effectiveness of our framework in managing heterogeneous multimodal EHR data, reducing false positives and potentially improving the allocation of medical resources in the ICU.
While this paper focuses on binary classification tasks, many clinical applications require solutions for multiclass tasks (e.g., disease diagnosis) and continuous regression tasks (e.g., survival prediction). These are equally important and relevant for advancing clinical practice. In the future, we plan to expand our framework by incorporating additional data modalities, such as time series and medical images, to provide deeper clinical insights and enhance model performance. We also aim to extend the framework to handle multimodal EHR multiclass tasks, offering valuable predictive guidance for complex clinical scenarios. Additionally, we intend to investigate regression tasks, leveraging the recently introduced Epistemic Random Fuzzy Set (ERFS) theory 32, 33 and further building on developments in evidential regression 34.
ACKNOWLEDGMENTS
This research is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG-GC-2019-001-2B).
APPENDIX
A  Dempster-Shafer (DS) Theory
Let $\Omega=\{\omega_{1},\omega_{2},\cdots,\omega_{M}\}$ be a finite set of hypotheses about some question, called the frame of discernment. Evidence about a variable taking values in $\Omega$ can be represented by a mass function: $2^{\Omega}$ to [0,1] such that $m(\emptyset)=0$ and
$\sum_{A\subseteq\Omega}m(A)=1.$ (1)
For any hypothesis $A\subseteq\Omega$, the quantity $m(A)$ is interpreted as a share of a unit mass of belief allocated to the hypothesis that the truth is in $A$, and which cannot be allocated to any strict subset of $A$ based on the available evidence. Set $A$ is called a focal set of $m$ if $m(A)>0$. A mass function is said to be Bayesian if its focal sets are singletons, and logical if it has only one focal set. Two mass functions $m_{1}$ and $m_{2}$ representing independent items of evidence can be combined conjunctively by Dempster’s combination rule 17 $\oplus$ as
$(m_{1}\oplus m_{2})(A)=\frac{\sum_{B\cap C=A}m_{1}(B)m_{2}(C)}{1-\sum_{B\cap C=\emptyset}m_{1}(B)m_{2}(C)},$ (2)
for all $A\neq\emptyset$, where $\sum_{B\cap C=\emptyset}m_{1}(B)m_{2}(C)$ is the degree of conflict among the two pieces of evidence, The nice information fusion attribute of DS theory points out the high potential of using DS theory for heterogenetic medical data analysis.
After aggregating all available evidence, we adopt the pignistic transformation, as proposed by Smets in the Transferable Belief Model 35, to make a final decision based on the combined mass function. This framework bases decisions on the pignistic probability distribution with the following expression:
$p(\omega)=\sum_{A\subseteq\Omega:\omega\in A}\frac{m(A)}{|A|},\forall\omega\in\Omega.$ (3)
B  Evidence mapping (EM)
As illustrated in Figure 6, the evidence mapping module consists of one input layer, one hidden layer, and one output layer. The input layer is composed of $H$ units ($H$ is the number of prototypes), whose weights vectors are prototypes $\pi_{1},\pi_{2},\cdots,\pi_{H}$ in input space. The activation of unit $h$ in the input layer is
$s_{h}=\beta_{h}\mathrm{exp}(-\gamma_{h}d_{h}^{2}),$ (4)
Figure 6: The illustration of evidence mapping
where $d_{h}=||x-\pi_{h}||$ denotes the Euclidean distance between input vector $x$ and prototype $\pi_{h}$ , $\gamma_{h}>0$ is a scale parameter, and $\beta_{h}\in[0,1]$ is an extra parameter. The hidden layer computes mass functions $m_{h}$ (evidence) of each prototype $\pi_{h}$ is defined as:
$m_{h}(\{\omega_{c}\})$ $=u_{h}^{(c)}s_{h},\quad c=1,2,\cdots,M,$ (5a)
$m_{h}(\Omega)$ $=1-s_{h},$ (5b)
where $u_{h}^{(c)}$ is the membership degree of prototype $h$ to class $\omega_{c}$, $\sum_{c=1}^{M}u_{h}^{(c)}=1$, and $M$ is the number of classes. Therefore, the vector of mass functions induced by prototypes is denoted as:
$m_{h}=(m_{h}(\{\omega_{1}\}),m_{h}(\{\omega_{2}\}),\cdots,m_{h}(\{\omega_{M}\}),m_{h}(\Omega))\in\mathbb{R}^{M+1}.$ (6)
Finally, the mass functions are then aggregated by Dempster’s combination rule using Eq. 2 in the output layer. A combined mass function $m$ is computed as the orthogonal sum of the $H$ mass functions:
$m=m_{1}\oplus m_{2}\oplus\cdots\oplus m_{H}\in\mathbb{R}^{M+1}.$ (7)
The combined mass functions (the outputs of the EM module) represent the degrees of belief about the given class with $m(\{\omega_{c}\})$, as well as its prediction uncertainty with $m(\Omega)$. In our binary classification case, the dimension of EM module outputs would be three.
C  Evidence fusion
To combine multiple mass functions $m_{1},m_{2},\cdots,m_{K}$ from different modalities/data types/data sources, Dempster’s combination rule is applied again Eq. 2 to aggregate evidence for multiple sources for final evidence generation.
$m=m_{1}\oplus m_{2}\oplus\cdots\oplus m_{K}\in\mathbb{R}^{M+1},$ (8)
where $K$ is the number of mass functions to combine. For example, $K$=2 for the fusion of evidence from structured EHRs and free-text notes.
D  The details of augmented learning algorithm
Let $p_{i}=(p_{i}(\omega_{1}),\cdots,p_{i}(\omega_{c}),\cdots,p_{i}(\omega_{M}))$ be the final probability after the pignistic transformation (3) for training sample $i$, and $y_{i}=(y_{i,1},y_{i,2},\cdots,y_{i,M})$ denotes the one-hot encoding for corresponding ground-truth labels. The main loss function $\mathcal{L}_{main}$ is computed as:
$\mathcal{L}_{p}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{c=1}^{M}w_{c}y_{i,c}\mathrm{log}(p_{i}(\omega_{c})),$ (9)
where $N$ is the number of training samples, $M$ denotes the number of classes, and $w_{c}$ is the weight assigned to each class to address the class imbalance issue.
Moreover, two auxiliary cross-entropy losses are introduced to optimize the feature representation performance of the encoders, as the evidence mapping module performs more effectively with high-quality representations. Firstly, to regulate the representation generated by encoders, we added an additional fully connected network (FCN) to generate logits $o_{i}$ for each modality. Let $o^{s}_{i}=(o_{i,1}^{s},o_{i,2}^{s},\cdots,o_{i,M}^{s})$ and $o^{n}_{i}=(o_{i,1}^{n},o_{i,2}^{n},\cdots,o_{i,M}^{n})$ be the logits from the encoders for structured data and free-text notes, respectively, the cross-entropy losses $\mathcal{L}^{s}_{aux}$ and $\mathcal{L}^{n}_{aux}$ are then calculated with $y_{i}$ for structured data and notes, respectively:
$\mathcal{L}^{s}_{aux}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{c=1}^{M}w_{c}y_{i,c}\mathrm{log}\frac{\mathrm{exp}(o_{i,c}^{s})}{\sum_{b=1}^{M}\mathrm{exp}(o_{i,b}^{s})},$ (10)
$\mathcal{L}^{n}_{aux}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{c=1}^{M}w_{c}y_{i,c}\mathrm{log}\frac{\mathrm{exp}(o_{i,c}^{n})}{\sum_{b=1}^{M}\mathrm{exp}(o_{i,b}^{n})},$ (11)
Ultimately, the overall loss function $\mathcal{L}_{overall}$ is defined as follows:
$\mathcal{L}_{overall}=\mathcal{L}_{main}+\alpha\mathcal{L}^{s}_{aux}+\beta\mathcal{L}^{n}_{aux},$ (12)
where $\alpha,\beta$ are the hyperparameters that control the balance between the main loss and the auxiliary cross-entropy losses.
E  The model comparison results in metrics (specificity, recall, sensitivity, and NPV)
[…]
Random Forest x 0.2634$\pm$0.0014 0.6946$\pm$0.0037 0.7398$\pm$0.0015 0.9476$\pm$0.0006
MLP x 0.2741$\pm$0.0023 0.7707$\pm$0.0045 0.7264$\pm$0.0047 0.9594$\pm$0.0005
ResNet x 0.2809$\pm$0.0007 0.7676$\pm$0.0036 0.7367$\pm$0.0017 0.9594$\pm$0.0005
FT-Transformer x 0.2838$\pm$0.0015 0.7832$\pm$0.0054 0.7352$\pm$0.0019 0.9620$\pm$0.0009
BERT x 0.1767$\pm$0.0030 0.6966$\pm$0.0225 0.5634$\pm$0.0217 0.9330$\pm$0.0021
BioBERT x 0.1793$\pm$0.0011 0.6029$\pm$0.0278 0.6299$\pm$0.0194 0.9225$\pm$0.0028
Clinical Bert x 0.1962$\pm$0.0012 0.7158$\pm$0.0080 0.6071$\pm$0.0073 0.9410$\pm$0.0009
Multimodal (BERT) x x 0.2761$\pm$0.0034 0.7807$\pm$0.0063 0.7255$\pm$0.0064 0.9611$\pm$0.0008
Multimodal (BioBERT) x x 0.2741$\pm$0.0031 0.7934$\pm$0.0065 0.7182$\pm$0.0067 0.9629$\pm$0.0008
Multimodal (Clinical Bert) x x 0.2908$\pm$0.0027 0.7678$\pm$0.0063 0.7489$\pm$0.0054 0.9601$\pm$0.0008
Ours (MLP-BERT) x x 0.3297$\pm$0.0076 0.6957$\pm$0.0168 0.8094$\pm$0.0114 0.9522$\pm$0.0020
Ours (MLP-BioBERT) x x 0.3377$\pm$0.0115 0.6792$\pm$0.0218 0.8191$\pm$0.0148 0.9503$\pm$0.0023
Ours (MLP-Clinical Bert) x x 0.3609$\pm$0.0075 0.6478$\pm$0.0123 0.8454$\pm$0.0079 0.9472$\pm$0.0013
Ours (ResNet-BERT) x x 0.3179$\pm$0.0073 0.7206$\pm$0.0168 0.7916$\pm$0.0118 0.9550$\pm$0.0020
Ours (ResNet-BioBERT) x x 0.3244$\pm$0.0100 0.7092$\pm$0.0219 0.8000$\pm$0.0150 0.9538$\pm$0.0025
Ours (ResNet-Clinical Bert) x x 0.326$\pm$0.0067 0.7151$\pm$0.0131 0.8010$\pm$0.0097 0.9546$\pm$0.0015
Ours (FT-Trans-BERT) x x 0.3163$\pm$0.0080 0.7433$\pm$0.0109 0.7836$\pm$0.0110 0.9580$\pm$0.0012
Ours (FT-Trans-BioBERT) x x 0.3282$\pm$0.0015 0.7193$\pm$0.0055 0.8027$\pm$0.0027 0.9553$\pm$0.0007
Ours (FT-Trans-Clinical Bert) x x 0.3289$\pm$0.0070 0.7371$\pm$0.0171 0.7974$\pm$0.0108 0.9578$\pm$0.0021
(a) Mortality prediction
[…]
Random Forest x 0.2360$\pm$0.0045 0.6744$\pm$0.0131 0.6616$\pm$0.0150 0.9295$\pm$0.0014
MLP x 0.2615$\pm$0.0013 0.6542$\pm$0.0034 0.7147$\pm$0.0030 0.9305$\pm$0.0004
ResNet x 0.2673$\pm$0.0016 0.6420$\pm$0.0039 0.7282$\pm$0.0025 0.9295$\pm$0.0006
FT-Transformer x 0.2603$\pm$0.0026 0.6974$\pm$0.0061 0.6938$\pm$0.0067 0.9369$\pm$0.0007
BERT x 0.2004$\pm$0.0031 0.5685$\pm$0.0147 0.6488$\pm$0.0160 0.9070$\pm$0.0009
BioBERT x 0.2006$\pm$0.0022 0.5504$\pm$0.0160 0.6610$\pm$0.0141 0.9051$\pm$0.0013
Clinical Bert x 0.2208$\pm$0.0024 0.6251$\pm$0.0133 0.6587$\pm$0.0116 0.9193$\pm$0.0014
Multimodal (BERT) x x 0.2568$\pm$0.0029 0.6892$\pm$0.0141 0.6915$\pm$0.0108 0.9347$\pm$0.0022
Multimodal (BioBERT) x x 0.2554$\pm$0.0034 0.6943$\pm$0.0136 0.6868$\pm$0.0114 0.9358$\pm$0.0018
Multimodal (Clinical Bert) x x 0.2507$\pm$0.0045 0.7201$\pm$0.0188 0.6664$\pm$0.0163 0.9394$\pm$0.0024
Ours (MLP-BERT) x x 0.2809$\pm$0.0038 0.6218$\pm$0.0050 0.7538$\pm$0.0064 0.9281$\pm$0.0005
Ours (MLP-BioBERT) x x 0.2892$\pm$0.0048 0.6022$\pm$0.0061 0.7710$\pm$0.0071 0.9262$\pm$0.0007
Ours (MLP-Clinical Bert) x x 0.2980$\pm$0.0075 0.5899$\pm$0.0134 0.7841$\pm$0.0117 0.9254$\pm$0.0012
Ours (ResNet-BERT) x x 0.2787$\pm$0.0071 0.6460$\pm$0.0158 0.7402$\pm$0.0148 0.9314$\pm$0.0016
Ours (ResNet-BioBERT) x x 0.2848$\pm$0.0059 0.6212$\pm$0.0126 0.7580$\pm$0.0119 0.9285$\pm$0.0012
Ours (ResNet-Clinical Bert) x x 0.2968$\pm$0.0029 0.6227$\pm$0.0073 0.7720$\pm$0.0058 0.9299$\pm$0.0008
Ours (FT-Trans-BERT) x x 0.2595$\pm$0.0036 0.7180$\pm$0.0116 0.6830$\pm$0.0110 0.9402$\pm$0.0014
Ours (FT-Trans-BioBERT) x x 0.2668$\pm$0.0042 0.6914$\pm$0.0106 0.7059$\pm$0.0109 0.9368$\pm$0.0012
Ours (FT-Trans-Clinical Bert) x x 0.2782$\pm$0.0026 0.6766$\pm$0.0046 0.7288$\pm$0.0050 0.9359$\pm$0.0005
(b) PLOS prediction
Table 3: The model comparison for mortality (a) and PLOS prediction (b) tasks in metrics (precision, recall, specificity, and NPV).
F  The evaluation on different fusion settings for PLOS prediction
Figure 7: The evaluation of our framework on different fusion settings for PLOS prediction: (1) modalities, (2) data types, (3) data sources.

### v2

@@
-Evidence-based multimodal fusion on structured EHRs and free-text notes for ICU outcome prediction
+Towards accurate and reliable ICU outcome prediction: a multimodal learning framework based on belief function theory using structured EHRs and free-text notes
 Abstract
-Objective: Accurate Intensive Care Unit (ICU) outcome prediction is critical for improving patient treatment quality and ICU resource allocation.
-Existing research mainly focuses on structured data and lacks effective frameworks to integrate clinical notes from heterogeneous electronic health records (EHRs).
-This study aims to explore a multimodal framework based on evidence theory that can effectively combine heterogeneous structured EHRs and free-text notes for accurate and reliable ICU outcome prediction.
-Materials and Methods: We proposed an evidence-based multimodal fusion framework to predict ICU outcomes, including mortality and prolonged length of stay (PLOS), by utilizing both structured EHR data and free-text notes from the MIMIC-III database.
-We compare the performance against baseline models that use only structured EHRs, free-text notes, or existing multimodal approaches.
-Results: The results demonstrate that the evidence-based multimodal fusion model achieved both accurate and reliable prediction.
+Accurate Intensive Care Unit (ICU) outcome prediction is critical for improving patient treatment quality and ICU resource allocation.
+Existing research mainly focuses on structured data, e.g.
+demographics and vital signs, and lacks effective frameworks to integrate clinical notes from heterogeneous electronic health records (EHRs).
+This study aims to explore a multimodal framework based on belief function theory that can effectively fuse heterogeneous structured EHRs and free-text notes for accurate and reliable ICU outcome prediction.
+The fusion strategy accounts for prediction uncertainty within each modality and conflicts between multimodal data.
+The experiments on MIMIC-III dataset show that our framework provides more accurate and reliable predictions than existing approaches.
 Specifically, it outperformed the best baseline by 1.05%/1.02% in BACC, 9.74%/6.04% in F1 score, 1.28%/0.9% in AUROC, and 6.21%/2.68% in AUPRC for predicting mortality and PLOS, respectively.
 Additionally, it improved the reliability of the predictions with a 26.8%/15.1% reduction in the Brier score and a 25.0%/13.3% reduction in negative log-likelihood.
-Conclusion: This study demonstrates that the evidence-based multimodal fusion framework can serve as a strong baseline for predictions using structured EHRs and free-text notes.
-It effectively reduces false positives, which can help improve the allocation of medical resources in the ICU.
-This framework can be further applied to analyze multimodal EHRs for other clinical tasks.
-Keywords: Multimodal learning, Evidence fusion, Dempster-Shafer theory, ICU outcome prediction, Electronic health records
-INTRODUCTION
+By effectively reducing false positives, the model can aid in better allocation of medical resources in the ICU.
+Furthermore, the proposed method is very versatile and can be extended to analyzing multimodal EHRs for other clinical tasks.
+The code implementation is available on https://github.com/yuchengruan/evid_multimodal_ehr.
+keywords:
+Multimodal Learning, Belief Function Theory, Evidence Fusion, ICU outcome prediction, Electronic Health Records
+1 Introduction
 The Intensive Care Unit (ICU) is a specialized hospital ward that offers comprehensive and continuous care to critically ill patients.
-As the population of critically ill patients grows, the demand for ICUs has risen significantly, placing strain on already limited and costly intensive care resources 1, especially during public health crises like the COVID-19 pandemic, when hospitals face an overwhelming surge of patients.
+As the population of critically ill patients grows, the demand for ICUs has risen significantly, placing strain on already limited and costly intensive care resources [1], especially during public health crises like the COVID-19 pandemic, when hospitals face an overwhelming surge of patients [2].
 Due to the limited availability of intensive care resources, researchers have emphasized the necessity of predicting ICU outcomes such as mortality rates and prolonged lengths of stay (PLOS).
 Accurate predictions can help in the efficient allocation of medical resources for patients in need and reduce unnecessary expenses without compromising patient care.
-Furthermore, they are crucial for healthcare providers in making informed decisions about patient care strategies and providing early interventions to patients at high risk of adverse outcomes 2, 3.
+Furthermore, they are crucial for healthcare providers in making informed decisions about patient care strategies and providing early interventions to patients at high risk of adverse outcomes [3, 4].
 Over the past two decades, the adoption of electronic ICU technology has enabled the collection of extensive data on ICU patients, creating new opportunities for developing advanced methods to predict ICU outcomes.
-Most previous research has concentrated on modeling ICU outcome predictions using structured EHR data 4, 5, 6, which often captures only a portion of clinical information.
+Most previous research has concentrated on modeling ICU outcome predictions using structured EHR data [5, 6, 7], which often captures only a portion of clinical information.
 It may miss out on the rich contextual information that unstructured EHR data (such as nursing notes, patient narratives, and imaging reports) can provide.
-Natural language processing (NLP) techniques have been well explored to extract valuable insights from unstructured free-text EHR data 7, 8, 9.
+Natural language processing (NLP) techniques have been well explored to extract valuable insights from unstructured free-text EHR data [8, 9, 10, 11].
 Therefore, effective multimodal learning algorithms are essential to integrate heterogeneous EHRs for better ICU outcome prediction.
-Recently, deep learning-based multimodal models have been proven to combine structured EHR data and free-text data at the deep feature level for clinical outcome predictions 10, 11, 12.
-While those methods have increased the prediction accuracy, the question regarding the reliability of those models in noisy and unstable clinical situations still exists.
-In this study, we hypothesize that an effective multimodal deep neural network that incorporates both structured EHR data and free-text notes can improve the prediction of ICU outcomes.
-Our objective is to propose and evaluate a novel evidence-based multimodal framework for ICU outcome prediction using multimodal EHR data.
-Unlike traditional methods that rely on data or feature embeddings, our framework integrates information based on evidence derived from different modalities.
+Recently, deep learning-based multimodal models have been proven to combine structured EHR data and free-text data at the deep feature level for clinical outcome predictions [12, 13, 14].
+These models often simply concatenate structured data with encoded features from free texts to generate patient representations for decision-making.
+While those approaches have improved prediction accuracy, the clinical impacts between the two heterogeneous modalities [15] are now well explored.
+Another limitation of existing research is the lack of reliability evaluation for deep learning models.
+Unreliable predictions can lead to incorrect diagnoses or treatment plans, potentially harming patients [16, 17, 18].
+Therefore, evaluating the prediction reliability is crucial beyond just predictive accuracy, especially in critical care settings.
+However, concerns about the reliability of existing models in noisy and unstable clinical environments still remain.
+Belief function theory (BFT), also known as Dempster-Shafer theory (DST), is a powerful framework for modeling, reasoning with, and integrating imperfect (noisy, uncertain, conflicting) data [19, 20, 21].
+The effectiveness of BFT in low-quality and multimodal medical image analysis has been widely reviewed in [22, 23].
+However, the study of BFT in EHR data is limited.
+Ling et al.[24] first studied the survival prediction uncertainty using structured EHRs under the framework of BFT and possibility theory.
+The heterogeneity of clinical data and other medical modalities, e.g., imaging and genetic, are also studied in [25] by combining multimodal data using the generated Dempster’s combination rule [26].
+In this work  [^1] This work is an extended version of the short paper presented at the 8th International Conference on Belief Functions (BELIEF 2024) [27]., we further study the effectiveness of BFT in multimodal EHR analysis using structured EHRs and free-text notes with a focus on ICU outcomes prediction.
+We propose a multimodal learning model under the BFT framework with accurate and reliable ICU outcomes prediction using multimodal EHR data.
+Instead of developing more effective feature extraction or interaction strategies for multimodal data communication, our framework focuses on effective evidence fusion study and integrates information based on the evidence derived from different modalities.
 Specifically, we use state-of-the-art deep neural networks for single-modality feature extraction: ResNet/Transformer-based models for structured EHR data and pre-trained language models for free-text EHR data.
-The extracted features are independently mapped into evidence and then combined in the evidence space.
-Experimental results on the MIMIC-III database for mortality and prolonged length of stay (PLOS) predictions demonstrate the effectiveness of the proposed model.
-The code implementation is available on GitHub repository (https://github.com/yuchengruan/evid_multimodal_ehr)
-MATERIALS AND METHODS
-Overview
-In this section, we demonstrate the cohort, patient features, prediction outcomes, and proposed evidence-based multimodal fusion framework in combining structured EHRs and free-text notes for ICU outcome prediction.
-Cohort
-MIMIC-III (Medical Information Mart for Intensive Care III) is a large, publicly available database containing de-identified health records from patients in critical care units at Beth Israel Deaconess Medical Center (from the US) between 2001 and 2012 13.
+The extracted features are independently mapped into evidence with an evidence mapping module and then combined in the evidence space in an evidence fusion module.
+Experimental results on the MIMIC-III database for mortality and prolonged length of stay (PLOS) predictions demonstrate the effectiveness of our proposed model in both predictive accuracy and reliability.
+2 Preliminaries
+2.1 Belief function theory
+Belief function theory (BFT) was first introduced by Dempster and Shafer [20, 19].
+The expressive capabilities of belief functions enable a more accurate representation of evidence than relying solely on probabilities.
+Let $\Omega=\{\omega_{1},\omega_{2},\cdots,\omega_{M}\}$ be a finite set of hypotheses about some question, called the frame of discernment.
+Evidence about a variable taking values in $\Omega$ can be represented by a mass function: $2^{\Omega}$ to [0,1] such that $m(\emptyset)=0$ and
+$\sum_{A\subseteq\Omega}m(A)=1.$ (1)
+For any hypothesis $A\subseteq\Omega$, the quantity $m(A)$ is interpreted as a share of a unit mass of belief allocated to the hypothesis that the truth is in $A$, and which cannot be allocated to any strict subset of $A$ based on the available evidence.
+Set $A$ is called a focal set of $m$ if $m(A)>0$.
+A mass function is said to be Bayesian if its focal sets are singletons, and logical if it has only one focal set.
+Two mass functions $m_{1}$ and $m_{2}$ representing independent items of evidence can be combined conjunctively by Dempster’s combination rule [19] $\oplus$ as
+$(m_{1}\oplus m_{2})(A)=\frac{\sum_{B\cap C=A}m_{1}(B)m_{2}(C)}{1-\sum_{B\cap C=\emptyset}m_{1}(B)m_{2}(C)},$ (2)
+for all $A\neq\emptyset$, where $\sum_{B\cap C=\emptyset}m_{1}(B)m_{2}(C)$ is the degree of conflict among the two pieces of evidence, The nice information fusion attribute of BFT points out the high potential in heterogenetic medical data analysis.
+After aggregating all available evidence, the final decision of BFT can be made based on the pignistic transformation proposed by Smets in the Transferable Belief Model [28] that combinese all mass functions using the following expression:
+$p(\omega)=\sum_{A\subseteq\Omega:\omega\in A}\frac{m(A)}{|A|},\forall\omega\in\Omega.$ (3)
+2.2 Evidential neural network
+Denœux [21] proposed an evidential neural network (ENN) that maps imperfect (uncertain, imprecise, or noise) input features into degrees of belief and ignorance (uncertainty) under the framework of BFT [19, 29].
+The essential concept of ENN is to consider each prototype as a piece of evidence, which is discounted based on its distance from the input vector.
+The evidence from different prototypes is then aggregated using Dempster’s combination rule.
+As illustrated in Figure 1, the ENN consists of one input layer, one hidden layer, and one output layer.
+The input layer is composed of $H$ units ($H$ is the number of prototypes), whose weights vectors are prototypes $\pi_{1},\pi_{2},\cdots,\pi_{H}$ in input space.
+The activation of unit $h$ in the input layer is
+$s_{h}=\beta_{h}\mathrm{exp}(-\gamma_{h}d_{h}^{2}),$ (4)
+Figure 1: The illustration of ENN
+where $d_{h}=||x-\pi_{h}||$ denotes the Euclidean distance between input vector $x$ and prototype $\pi_{h}$ , $\gamma_{h}>0$ is a scale parameter, and $\beta_{h}\in[0,1]$ is an extra parameter.
+The hidden layer computes mass functions $m_{h}$ (evidence) of each prototype $\pi_{h}$ is defined as:
+$m_{h}(\{\omega_{c}\})$ $=u_{h}^{(c)}s_{h},\quad c=1,2,\cdots,M,$ (5a)
+$m_{h}(\Omega)$ $=1-s_{h},$ (5b)
+where $u_{h}^{(c)}$ is the membership degree of prototype $h$ to class $\omega_{c}$, $\sum_{c=1}^{M}u_{h}^{(c)}=1$, and $M$ is the number of classes.
+Therefore, the vector of mass functions induced by prototypes is denoted as:
+$m_{h}=(m_{h}(\{\omega_{1}\}),m_{h}(\{\omega_{2}\}),\cdots,m_{h}(\{\omega_{M}\}),m_{h}(\Omega))\in\mathbb{R}^{M+1}.$ (6)
+Finally, the mass functions are then aggregated by Dempster’s combination rule using Eq.
+2 in the output layer.
+A combined mass function $m$ is computed as the orthogonal sum of the $H$ mass functions:
+$m=m_{1}\oplus m_{2}\oplus\cdots\oplus m_{H}\in\mathbb{R}^{M+1}.$ (7)
+The combined mass functions (the outputs of the ENN) represent the degrees of belief about the given class with $m(\{\omega_{c}\})$, as well as its prediction uncertainty with $m(\Omega)$.
+In our binary classification case, the dimension of ENN outputs would be three.
+3 METHODS
+3.1 Model architecture
+The overview of the proposed framework is illustrated in Figure 2.
+The general idea of this framework is to generate the modality-level evidence for both structured data and free-text notes using the evidence mapping module and fuse the modality-level evidence with Dempster’s combination rule for final prediction.
+Figure 2: The overview of our proposed framework. EM: evidence mapping and EF: evidence fusion
+3.1.1 Evidence mapping (EM)
+Inspired by ENN and its promising adoptions in medical data analysis [30, 31, 32, 33], we propose incorporating ENN as an evidence mapping module with the state-of-the-art encoders to generate evidence for structured EHRs and free-text notes.
+Given modality level input, the evidence mapping module can output the evidence for each class as well as the uncertainty regarding this prediction.
+Structured data evidence mapping
+To produce modality evidence for structured data, we initially used a structured data encoder to extract deep features (the output dimension is set to 32).
+We considered three popular encoders to extract the embeddings: MLP, ResNet, FT-Transformer [34] (see Baselines section for more details).
+Subsequently, we introduced an evidence mapping module (the number of prototypes is set to 20) to transform the deep features into evidence embeddings for structured data.
+Free-text notes evidence mapping
+Similarly, we utilized pre-trained language models to extract the deep features from clinical notes, on which we developed an evidence mapping module (the number of prototypes is set to 20) to produce the evidence of the modality.
+Our primary analysis focused on pre-trained architectures similar to BERT.
+Accordingly, three BERT-based architectures were evaluated: BERT [35], BioBERT [35], Clinical BERT [36] (see Baselines section for more details).
+To minimize computational overhead while maintaining high predictive performance, we froze the pre-trained language models and fine-tuned an additional layer (128 hidden units) on top in model training.
+3.1.2 Evidence fusion (EF)
+Based on the modality-level evidence obtained from structured EHR and free-text notes, we developed the evidence fusion module based on Dempster’s combination rule to generate the final evidence for decision-making.
+To combine multiple mass functions $m_{1},m_{2},\cdots,m_{K}$ from different modalities/data types/data sources, Dempster’s combination rule is applied again Eq.
+2 to aggregate evidence for multiple sources for final evidence generation.
+$m=m_{1}\oplus m_{2}\oplus\cdots\oplus m_{K}\in\mathbb{R}^{M+1},$ (8)
+where $K$ is the number of mass functions to combine. For example, $K$=2 for the fusion of evidence from structured EHRs and free-text notes.
+3.2 Augmented model optimization algorithm
+We optimize the proposed framework using an augmented learning algorithm, which includes two types of optimization objectives: (1) main objective and (2) auxiliary objective.
+The main objective is to optimize predictive performance based on transformed evidence as the primary loss function.
+Additionally, two auxiliary cross-entropy losses are incorporated to enhance the feature representation capability of the independent encoders for the two modalities, as the evidence mapping module performs more effectively with high-quality representations.
+Let $p_{i}=(p_{i}(\omega_{1}),\cdots,p_{i}(\omega_{c}),\cdots,p_{i}(\omega_{M}))$ be the final probability after the pignistic transformation (3) for training sample $i$, and $y_{i}=(y_{i,1},y_{i,2},\cdots,y_{i,M})$ denotes the one-hot encoding for corresponding ground-truth labels.
+The main loss function $\mathcal{L}_{main}$ is computed as:
+$\mathcal{L}_{main}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{c=1}^{M}w_{c}y_{i,c}\mathrm{log}(p_{i}(\omega_{c})),$ (9)
+where $N$ is the number of training samples, $M$ denotes the number of classes, and $w_{c}$ is the weight assigned to each class to address the class imbalance issue.
+Moreover, two auxiliary cross-entropy losses are introduced to optimize the feature representation performance of the encoders, as the evidence mapping module performs more effectively with high-quality representations.
+Firstly, to regulate the representation generated by encoders, we added an additional fully connected network (FCN) to generate logits $o_{i}$ for each modality.
+Let $o^{s}_{i}=(o_{i,1}^{s},o_{i,2}^{s},\cdots,o_{i,M}^{s})$ and $o^{n}_{i}=(o_{i,1}^{n},o_{i,2}^{n},\cdots,o_{i,M}^{n})$ be the logits from the encoders for structured data and free-text notes, respectively, the cross-entropy losses $\mathcal{L}^{s}_{aux}$ and $\mathcal{L}^{n}_{aux}$ are then calculated with $y_{i}$ for structured data and notes, respectively:
+$\mathcal{L}^{s}_{aux}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{c=1}^{M}w_{c}y_{i,c}\mathrm{log}\frac{\mathrm{exp}(o_{i,c}^{s})}{\sum_{b=1}^{M}\mathrm{exp}(o_{i,b}^{s})},$ (10)
+$\mathcal{L}^{n}_{aux}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{c=1}^{M}w_{c}y_{i,c}\mathrm{log}\frac{\mathrm{exp}(o_{i,c}^{n})}{\sum_{b=1}^{M}\mathrm{exp}(o_{i,b}^{n})},$ (11)
+Ultimately, the overall loss function $\mathcal{L}_{overall}$ is defined as follows:
+$\mathcal{L}_{overall}=\mathcal{L}_{main}+\alpha\mathcal{L}^{s}_{aux}+\beta\mathcal{L}^{n}_{aux},$ (12)
+where $\alpha,\beta$ are the hyperparameters that control the balance between the main loss and the auxiliary cross-entropy losses. In both tasks, we set $\alpha=2$ and $\beta=1$.
+4 Experiments
+4.1 Study Cohort
+This study used data from MIMIC-III (Medical Information Mart for Intensive Care III), a large, publicly available database containing de-identified health records from patients in critical care units at Beth Israel Deaconess Medical Center (from the US) between 2001 and 2012 [37].
 We collected structured EHR data and free-text clinical notes from the database.
@@
 For patients with multiple ICU stays, we considered only the first.
-Patient features
-Patient features include both structured EHR data and unstructured free-text notes.
+4.2 Input features
+Input features include both structured EHR data and unstructured free-text notes.
 In this section, we demonstrate the patient features in our study and provide details about the data preprocessing steps.
@@
 As a result, NLP techniques, especially pre-trained LLM, can be applied to these notes to gain deeper insights for data-driven predictions.
-In this study, we focus on Nursing, Nursing/Other, Physician, and Radiology notes, as these comprise the majority of clinical documentation and are frequently recorded in the MIMIC-III database 10.
+In this study, we focus on Nursing, Nursing/Other, Physician, and Radiology notes, as these comprise the majority of clinical documentation and are frequently recorded in the MIMIC-III database [12].
 We extracted only the first 24 hours of notes for each admission to facilitate early outcome prediction.
 […]
-Prediction Outcomes
+4.3 Prediction Tasks
 In this study, we focus on two ICU prediction outcomes: mortality and prolonged length of stay.
+Since the two clinical outcomes are rare in patient popularity, we applied a simple class weighting approach during training based on relative class frequencies to mitigate biases and handle the imbalance in EHR data.
 Mortality
@@
 Length of stay refers to the duration between a patient’s admission to and discharge from the ICU.
-In this study, we aim to predict prolonged length of stay (PLOS), defined as a stay exceeding 7 days 14, 15, 10.
+In this study, we aim to predict prolonged length of stay (PLOS), defined as a stay exceeding 7 days [38, 39, 12].
 Prolonged ICU stays are often linked to severe illnesses, complications, and increased mortality.
@@
 This problem is framed as a binary classification task.
-Model architecture
-The overview of the proposed framework is illustrated in Figure 1.
-In general, we propose to generate the modality-level evidence for both structured data and free-text notes using the evidence mapping module and fuse the modality-level evidence for final prediction.
-Figure 1: The overview of our proposed framework. EM: evidence mapping and EF: evidence fusion
-Evidence mapping (EM)
-Denœux 16 proposed an evidential neural network (ENN) that maps imperfect (uncertain, imprecise, or noise) input features into degrees of belief and ignorance (uncertainty) under the framework of Dempster-Shafer (DS) theory 17, 18.
-The essential concept of ENN is to consider each prototype as a piece of evidence, which is discounted based on its distance from the input vector.
-The evidence from different prototypes is then aggregated using Dempster’s combination rule.
-Inspired by ENN and its promising adoptions in medical data analysis 19, 20, 21, 22, we propose incorporating ENN as an evidence mapping module with the state-of-the-art encoders to generate evidence for structured EHRs and free-text notes.
-Further details about DS theory and evidence mapping can be found in Appendix A and B.
-Given modality level input, the evidence mapping module can output the evidence for each class as well as the uncertainty regarding this prediction.
-Structured data evidence mapping
-To produce modality evidence for structured data, we initially used a structured data encoder to extract deep features (the output dimension is set to 32).
-We considered three popular encoders to extract the embeddings: MLP, ResNet, FT-Transformer 23 (see Baselines section for more details).
-Subsequently, we introduced an evidence mapping module to transform the deep features into evidence embeddings for structured data.
-Free-text notes evidence mapping
-Similarly, we utilized pre-trained language models to extract the deep features from clinical notes, on which we developed an evidence mapping module to produce the evidence of the modality.
-Our primary analysis focused on pre-trained architectures similar to BERT.
-Accordingly, three BERT-based architectures were evaluated: BERT 24, BioBERT 24, Clinical BERT 25 (see Baselines section for more details).
-To minimize computational overhead while maintaining high predictive performance, we froze the pre-trained language models and fine-tuned an additional layer (128 hidden units) on top in model training.
-Evidence fusion (EF)
-Based on the modality-level evidence obtained from structured EHR and free-text notes, we developed the evidence fusion module based on Dempster’s combination rule to generate the final evidence for decision-making.
-Further details about the fusion of EHR and free-text notes evidence can be found in Appendix C.
-Augmented learning algorithm
-We optimize the proposed framework using an augmented learning algorithm, which includes two types of optimization objectives: (1) main objective and (2) auxiliary objective.
-The main objective is to optimize predictive performance based on transformed evidence as the primary loss function.
-Additionally, two auxiliary cross-entropy losses are incorporated to enhance the feature representation capability of the independent encoders for the two modalities, as the evidence mapping module performs more effectively with high-quality representations.
-More details can be found in the Appendix D.
-Evaluation and experiment
-Baselines
-To comprehensively evaluate the effectiveness of our proposed fusion framework, we employed three types of classification models as baselines: (1) models with structured data only, (2) models with free-text notes only, and (3) existing multimodal models based on both data.
-The configurations for the structured data models are as follows:
-Random Forest 26: A decision tree-based ensemble learning method for making predictions.
-MLP 23: A fundamental neural network with fully connected layers to encode structured data and serves as a reliable baseline.
+4.4 Baselines
+To comprehensively evaluate the effectiveness of our proposed fusion framework, we compared it against three baseline model categories: (1) models using only structured data, (2) models using only free-text notes, and (3) existing multimodal models that integrate both data types.
+Structured data baseline
+The following models were used to evaluate performance with structured EHR data:
+  • Random Forest [40]: A decision tree-based ensemble learning method for making predictions.
+  • MLP [34]: A fundamental neural network with fully connected layers to encode structured data and serves as a reliable baseline.
 The model was configured with 3 layers, 32 hidden units, and a dropout rate of 0.1.
-ResNet 23: Because of the success of ResNet in computer vision 27, it has also been adapted for structured data modeling.
+  • ResNet [34]: Because of the success of ResNet in computer vision [41], it has also been adapted for structured data modeling.
 Specifically, the main building block is simplified by providing a direct path from input to output.
 The configuration in our study has 3 residual blocks, 32 hidden units, and a dropout rate of 0.1.
-FT-Transformer 23: It converts all categorical and numerical features into embeddings, which are then processed through a couple of Transformer layers.
+  • FT-Transformer [34]: It converts all categorical and numerical features into embeddings, which are then processed through a couple of Transformer layers.
 It has demonstrated superior performance as a structured data encoder across various tasks.
 The model configuration has 3 Transformer layers with 192 hidden units, 8 attention heads, and a dropout rate of 0.2.
-As our focus is on pre-trained language models similar to BERT, we employed three BERT-based text classification models in this study:
-BERT 24: a pre-trained large language model English corpus using a self-supervised approach.
-Through masked language modeling and next-sentence prediction objectives, the model learns internal representations of the English language, enabling it to extract features useful for downstream tasks.
-In this study, we utilized the google-bert/bert-base-uncased model from Huggingface 28 for feature extraction.
-BioBERT 29: a variant of BERT pre-trained on biomedical literature, such as PubMed abstracts, and is optimized to perform more effectively on biomedical NLP tasks.
+Free-text notes baseline
+To assess text-based prediction performance, we compared our model against three BERT-based text classification approaches:
+  • BERT [35]: A pre-trained language model trained on a large English corpus using self-supervised learning.
+It learns contextual representations through masked language modeling and next-sentence prediction.
+In this study, we used the Google-bert/bert-base-uncased model from Huggingface [42] for feature extraction.
+  • BioBERT [43]: a variant of BERT pre-trained on biomedical literature, such as PubMed abstracts, and is optimized to perform more effectively on biomedical NLP tasks.
 In this study, we used dmis-lab/biobert-v1.1 from Huggingface as the extraction model.
-Clinical BERT 25: A fine-tuned version of BERT on clinical notes from the MIMIC-III database, making it well-suited for handling medical terminology and clinical narratives to enhance performance on clinical tasks.
+  • Clinical BERT [36]: A fine-tuned version of BERT on clinical notes from the MIMIC-III database, making it well-suited for handling medical terminology and clinical narratives to enhance performance on clinical tasks.
 In our study, we used emilyalsentzer/Bio_ClinicalBERT from the Huggingface transformer library for extracting embeddings from clinical notes.
-We also implemented a Multimodal approach for the ICU outcome prediction from 12: it concatenated the structured EHRs with the extracted embeddings from free-text notes, which were then passed into 2 fully connected layers for prediction.
-It supports different text encoders, and in this study, we used BERT, BioBERT, and Clinical BERT for fair comparisons
-Implementation details
+Multimodal modal baseline
+To compare against multimodal models, we implemented the concatenation-based approach from [14].
+This method combines structured EHR data with extracted text embeddings from free-text notes, followed by two fully connected layers for prediction.
+To ensure fairness, we tested this approach with BERT, BioBERT, and Clinical BERT as text encoders.
+4.5 Implementation details
 The dataset was randomly divided with 60% for training, 20% for validation, and 20% for testing.
 For model training, a mini-batch size of 32 was used, and the maximum number of epochs was set to 150, with early stopping applied.
+To handle data imbalance, we used a simple class weighting technique  [^2] https://scikit-learn.org/stable/modules/generated/sklearn.utils.class_weight.compute_class_weight.html based on class frequencies, as this was not the main focus of our research.
+The positive to negative weight ratios were set to 4.254:0.567 for the mortality prediction task and 3.660:0.579 for the PLOS prediction task.
 Each model was trained five times with different random seeds, and we reported the average results along with the standard error of the metrics to ensure statistical reliability.
 Hyperparameters were optimized for all baseline models to achieve the best results.
-All compared models were implemented using Scikit-learn 30, PyTorch 31, and Hugging Face’s Transformers library 28 in Python 3.8.19.
-The MLP, ResNet, and FT-Transformer models were built using the original source code on GitHub (https://github.com/yandex-research/rtdl-revisiting-models).
-The Multimodal approach was modified based on the code on Github (https://github.com/WeiChunLin/Bio_Clinical_BERT_Multimodal_Model).
-Model evaluation
-For comprehensive model evaluation and comparison, we reported the following metrics: precision, recall, specificity, negative predictive value (NPV), balanced accuracy (BACC), F1 score, area under the receiver operating characteristic curve (AUROC), area under the precision-recall curve (AUPRC), Brier score, and negative log-likelihood (NLL).
-Precision, recall, specificity, and NPV assess model performance for positive or negative samples.
-In contrast, BACC, F1 score, AUROC, and AUPRC provide a more holistic evaluation of predictive performance.
-Meanwhile, the Brier score and NLL measure the prediction reliability.
-Other fusion settings
-In addition to fusing evidence from modalities (structured EHRs and free-text notes), we evaluated our framework under two additional fusion settings: data types and data sources.
-The illustration is shown in Figure 2.
-Figure 2: The illustration of different fusion settings.
-Data types: Features can be grouped based on their data types. In our study, we further divided the structured data into two types: numerical and categorical data.
-Data Sources: Features can also be grouped according to their sources.
-In our case, the structured data was split into four sources: demographics, vital signs/lab tests, medical treatments, and comorbidities.
-Similarly, we categorized the clinical notes into four types: Nursing, Nursing/Other, Physician, and Radiology notes.
-For each group, we generated evidence using individual encoders and evidential layers.
-These pieces of evidence were then aggregated through the EF module to produce the final evidence for decision-making.
-RESULTS
-Data description
+All compared models were implemented using Scikit-learn [44], PyTorch [45], and Hugging Face’s Transformers library [42] in Python 3.8.19.
+The MLP, ResNet, and FT-Transformer models were built using the original source code on GitHub  [^3] https://github.com/yandex-research/rtdl-revisiting-models..
+The Multimodal approach was modified based on the code on Github  [^4] https://github.com/WeiChunLin/Bio_Clinical_BERT_Multimodal_Model..
+4.6 Model evaluation
+For comprehensive model evaluation and comparison, we reported the two types of metrics: predictive accuracy and reliability.
+  • Predictive accuracy ensures that models correctly identify critically ill patients who require urgent intervention, thereby reducing the risk of misdiagnosis and unnecessary treatments.
+To comprehensively assess accuracy, we consider two aspects:
+    • Class-specific accuracy metrics which evaluate performance separately for positive and negative cases using metrics such as precision, recall, specificity, and negative predictive value (NPV).
+      Precision $=\frac{\text{TP}}{\text{TP}+\text{FP}},$ (13)
+      Recall $=\frac{\text{TP}}{\text{TP}+\text{FN}},$ (14)
+      Specificity $=\frac{\text{TN}}{\text{TN}+\text{FP}},$ (15)
+      NPV $=\frac{\text{TN}}{\text{TN}+\text{FN}},$ (16)
+      where TP, TN, FP, and FN are True Positive, True Negative, False Positive, and False Negative, respectively.
+    • Holistic accuracy metrics include balanced accuracy (BACC), F1 score, the area under the receiver operating characteristic curve (AUROC), and the area under the precision-recall curve (AUPRC).
+AUROC is determined by calculating the area under the ROC curve (TP rate against FP rate across different threshold settings) and AUPRC calculates the area under the Precision-Recall curve across various threshold settings.
+      BACC $=\frac{1}{2}\left(\text{Recall}+\text{Specificity}\right),$ (17)
+      F1 $=2\cdot\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}.$ (18)
+      BACC, F1 score, and AUPRC are well-suited for evaluating model performance in imbalanced class distributions, as they provide a more balanced assessment than traditional accuracy measures.
+  • Reliability metrics quantify the model’s confidence in predictions.
+A model with high reliability provides well-calibrated probability estimates, allowing clinicians to make informed risk assessments.
+Here, we evaluate the performance of the Brier score and negative log-likelihood (NLL).
+    Brier Score $=\frac{1}{N}\sum_{i=1}^{N}(p_{i}-o_{i})^{2},$ (19)
+    NLL $=-\frac{1}{N}\sum_{i=1}^{N}\left(o_{i}\log(p_{i})+(1-o_{i})\log(1-p_{i})\right),$ (20)
+    where $N$ is the number of instances, $p_{i}$ is the predicted probability of the positive class for instance $i$, and $o_{i}$ is the actual outcome for instance $i$ (1 if positive, 0 if negative).
+5 RESULTS
+5.1 Data description
 Table 1 shows the descriptive characteristics of the patients in the study cohort.
@@
 Over half (51.2%) of patients received mechanical ventilation during their ICU stay and hypertension (38.8%) and congestive heart failure (31.2%) were among the most common patient comorbidities.
-Patient chracteristic Mortality PLOS
+Patient characteristic Mortality PLOS
 Yes (N=4540) No (N=33928) Yes (N=5220) No (N=33248)
@@
   Hispanic 88 (1.9%) 803 (2.4%) 104 (2.0%) 805 (2.4%)
-  Other/Unknown 986 (21.7%) 4928 (14.5%) 948 (18.2%) 4966 (14.9%)
+  Other 986 (21.7%) 4928 (14.5%) 948 (18.2%) 4966 (14.9%)
 Admission type
@@
 For continuous features, the mean and standard deviation are reported in the study.
-Model performance
-Table 2 presents the comparison of model performance under predictive evaluation metrics (BACC, F1 score, AUROC, and AUPRC) and reliability evaluation metrics (Brier score and NLL) for both mortality prediction (2(a)) and PLOS (2(b)) prediction tasks.
-The comparison results under precision, recall, specificity, and NPV can be found in Appendix E.
+5.2 Model performance
+5.2.1 Predictive accuracy
+Table 2 and 3 present the comparison of model performance for holistic predictive accuracy in mortality and PLOS prediction tasks, respectively.
 […]
-Random Forest x 0.7172$\pm$0.0019 0.3820$\pm$0.0018 0.7988$\pm$0.0010 0.3766$\pm$0.0030 0.1891$\pm$0.0006 0.5639$\pm$0.0015
-MLP x 0.7486$\pm$0.0003 0.4043$\pm$0.0019 0.8326$\pm$0.0011 0.4429$\pm$0.0033 0.1677$\pm$0.0015 0.4937$\pm$0.0037
-ResNet x 0.7521$\pm$0.0011 0.4113$\pm$0.0007 0.8350$\pm$0.0006 0.4468$\pm$0.0016 0.1654$\pm$0.0006 0.4895$\pm$0.0015
-FT-Transformer x 0.7592$\pm$0.0025 0.4166$\pm$0.0021 0.8426$\pm$0.0013 0.4577$\pm$0.0029 0.1648$\pm$0.0014 0.4832$\pm$0.0037
-BERT x 0.6300$\pm$0.0014 0.2814$\pm$0.0019 0.6777$\pm$0.0014 0.2037$\pm$0.0012 0.2302$\pm$0.0070 0.6523$\pm$0.0159
-BioBERT x 0.6164$\pm$0.0043 0.2759$\pm$0.0021 0.6695$\pm$0.0016 0.2181$\pm$0.0009 0.2214$\pm$0.0021 0.6326$\pm$0.0048
-Clinical Bert x 0.6614$\pm$0.0007 0.3080$\pm$0.0009 0.7240$\pm$0.0001 0.2928$\pm$0.0008 0.2132$\pm$0.0030 0.6086$\pm$0.0070
-Multimodal (BERT) x x 0.7531$\pm$0.0016 0.4079$\pm$0.0031 0.8398$\pm$0.0006 0.4596$\pm$0.0021 0.1691$\pm$0.0030 0.4992$\pm$0.0077
-Multimodal (BioBERT) x x 0.7558$\pm$0.0007 0.4073$\pm$0.0027 0.8362$\pm$0.0006 0.4570$\pm$0.0032 0.1725$\pm$0.0028 0.5084$\pm$0.0088
-Multimodal (Clinical Bert) x x 0.7584$\pm$0.0005 0.4218$\pm$0.0019 0.8404$\pm$0.0004 0.4686$\pm$0.0026 0.1606$\pm$0.0022 0.4792$\pm$0.0046
-Ours (MLP-BERT) x x 0.7525$\pm$0.0030 0.4466$\pm$0.0042 0.8406$\pm$0.0010 0.4764$\pm$0.0020 0.1345$\pm$0.0050 0.4041$\pm$0.0128
-Ours (MLP-BioBERT) x x 0.7492$\pm$0.0042 0.4493$\pm$0.0060 0.8408$\pm$0.0012 0.4758$\pm$0.0009 0.1301$\pm$0.0062 0.3927$\pm$0.0159
-Ours (MLP-Clinical Bert) x x 0.7467$\pm$0.0022 0.4629$\pm$0.0033 0.8465$\pm$0.0008 0.4935$\pm$0.0011 0.1176$\pm$0.0027 0.3594$\pm$0.0066
-Ours (ResNet-BERT) x x 0.7581$\pm$0.0037 0.4404$\pm$0.0038 0.8415$\pm$0.0011 0.4688$\pm$0.0014 0.1435$\pm$0.0052 0.4296$\pm$0.0133
-Ours (ResNet-BioBERT) x x 0.7546$\pm$0.0039 0.4438$\pm$0.0053 0.8424$\pm$0.0012 0.4671$\pm$0.0014 0.1392$\pm$0.0061 0.4187$\pm$0.0157
-Ours (ResNet-Clinical Bert) x x 0.7580$\pm$0.0023 0.4472$\pm$0.0042 0.8474$\pm$0.0006 0.4899$\pm$0.0015 0.1387$\pm$0.0043 0.4178$\pm$0.0110
-Ours (FT-Trans-BERT) x x 0.7634$\pm$0.0012 0.4432$\pm$0.0060 0.8485$\pm$0.0007 0.4797$\pm$0.0029 0.1450$\pm$0.0049 0.4345$\pm$0.0129
-Ours (FT-Trans-BioBERT) x x 0.7610$\pm$0.0015 0.4507$\pm$0.0007 0.8486$\pm$0.0011 0.4796$\pm$0.0027 0.1358$\pm$0.0014 0.4095$\pm$0.0045
-Ours (FT-Trans-Clinical Bert) x x 0.7672$\pm$0.0032 0.4541$\pm$0.0032 0.8534$\pm$0.0012 0.4977$\pm$0.0011 0.1390$\pm$0.0043 0.4192$\pm$0.0112
-(a) Mortality prediction
+Random Forest x 0.7172$\pm$0.0019 0.3820$\pm$0.0018 0.7988$\pm$0.0010 0.3766$\pm$0.0030
+MLP x 0.7486$\pm$0.0003 0.4043$\pm$0.0019 0.8326$\pm$0.0011 0.4429$\pm$0.0033
+ResNet x 0.7521$\pm$0.0011 0.4113$\pm$0.0007 0.8350$\pm$0.0006 0.4468$\pm$0.0016
+FT-Transformer x 0.7592$\pm$0.0025 0.4166$\pm$0.0021 0.8426$\pm$0.0013 0.4577$\pm$0.0029
+BERT as text encoder
+Text encoder only x 0.6300$\pm$0.0014 0.2814$\pm$0.0019 0.6777$\pm$0.0014 0.2037$\pm$0.0012
+Multimodal x x 0.7531$\pm$0.0016 0.4079$\pm$0.0031 0.8398$\pm$0.0006 0.4596$\pm$0.0021
 […]
-Random Forest x 0.6680$\pm$0.0025 0.3492$\pm$0.0038 0.7270$\pm$0.0018 0.2980$\pm$0.0029 0.2350$\pm$0.0009 0.6649$\pm$0.0018
-MLP x 0.6845$\pm$0.0009 0.3736$\pm$0.0011 0.7528$\pm$0.0010 0.3353$\pm$0.0010 0.1950$\pm$0.0018 0.5739$\pm$0.0044
-ResNet x 0.6851$\pm$0.0017 0.3774$\pm$0.0018 0.7532$\pm$0.0004 0.3366$\pm$0.0021 0.1929$\pm$0.0012 0.5702$\pm$0.0027
-FT-Transformer x 0.6956$\pm$0.0011 0.3790$\pm$0.0021 0.7674$\pm$0.0006 0.3404$\pm$0.0021 0.1947$\pm$0.0031 0.5704$\pm$0.0079
-BERT x 0.6087$\pm$0.0011 0.2960$\pm$0.0016 0.6507$\pm$0.0015 0.2262$\pm$0.0009 0.2316$\pm$0.0043 0.6557$\pm$0.0091
-BioBERT x 0.6057$\pm$0.0016 0.2940$\pm$0.0012 0.6487$\pm$0.0014 0.2214$\pm$0.0012 0.2297$\pm$0.0029 0.6518$\pm$0.0063
-Clinical Bert x 0.6420$\pm$0.0016 0.3261$\pm$0.0015 0.6946$\pm$0.0009 0.2661$\pm$0.0006 0.2225$\pm$0.0031 0.6350$\pm$0.0069
-Multimodal (BERT) x x 0.6903$\pm$0.0018 0.3739$\pm$0.0012 0.7625$\pm$0.0006 0.3427$\pm$0.0007 0.2017$\pm$0.0028 0.5960$\pm$0.0082
-Multimodal (BioBERT) x x 0.6905$\pm$0.0022 0.3731$\pm$0.0025 0.7606$\pm$0.0007 0.3364$\pm$0.0009 0.1974$\pm$0.0026 0.5888$\pm$0.0043
-Multimodal (Clinical Bert) x x 0.6933$\pm$0.0014 0.3714$\pm$0.0023 0.7648$\pm$0.0009 0.3544$\pm$0.0018 0.2027$\pm$0.0056 0.5871$\pm$0.0136
-Ours (MLP-BERT) x x 0.6878$\pm$0.0016 0.3868$\pm$0.0030 0.7580$\pm$0.0004 0.3499$\pm$0.0015 0.1773$\pm$0.0022 0.5288$\pm$0.0063
-Ours (MLP-BioBERT) x x 0.6866$\pm$0.0022 0.3905$\pm$0.0038 0.7573$\pm$0.0010 0.3461$\pm$0.0023 0.1720$\pm$0.0031 0.5172$\pm$0.0076
-Ours (MLP-Clinical Bert) x x 0.6870$\pm$0.0012 0.3952$\pm$0.0032 0.7633$\pm$0.0008 0.3557$\pm$0.0021 0.1637$\pm$0.0054 0.4943$\pm$0.0145
-Ours (ResNet-BERT) x x 0.6931$\pm$0.0013 0.3887$\pm$0.0040 0.7655$\pm$0.0011 0.3584$\pm$0.0007 0.1815$\pm$0.0045 0.5398$\pm$0.0107
-Ours (ResNet-BioBERT) x x 0.6896$\pm$0.0011 0.3918$\pm$0.0037 0.7647$\pm$0.0009 0.3532$\pm$0.0010 0.1752$\pm$0.0048 0.5241$\pm$0.0123
-Ours (ResNet-Clinical Bert) x x 0.6974$\pm$0.0011 0.4019$\pm$0.0015 0.7705$\pm$0.0009 0.3639$\pm$0.0010 0.1694$\pm$0.0018 0.5082$\pm$0.0042
-Ours (FT-Trans-BERT) x x 0.7005$\pm$0.0010 0.3809$\pm$0.0024 0.7725$\pm$0.0006 0.3524$\pm$0.0011 0.1927$\pm$0.0039 0.5617$\pm$0.0100
-Ours (FT-Trans-BioBERT) x x 0.6986$\pm$0.0011 0.3847$\pm$0.0030 0.7701$\pm$0.0012 0.3504$\pm$0.0014 0.1832$\pm$0.0045 0.5376$\pm$0.0113
-Ours (FT-Trans-Clinical Bert) x x 0.7027$\pm$0.0011 0.3942$\pm$0.0020 0.7743$\pm$0.0006 0.3575$\pm$0.0014 0.1760$\pm$0.0018 0.5205$\pm$0.0038
-(b) PLOS prediction
-Table 2: The model comparison for mortality (a) and PLOS prediction (b) tasks.
+Ours (ResNet) x x 0.7581$\pm$0.0037 0.4404$\pm$0.0038 0.8415$\pm$0.0011 0.4688$\pm$0.0014
+Ours (FT-Trans) x x 0.7634$\pm$0.0012 0.4432$\pm$0.0060 0.8485$\pm$0.0007 0.4797$\pm$0.0029
+BioBERT as text encoder
+Text encoder only x 0.6164$\pm$0.0043 0.2759$\pm$0.0021 0.6695$\pm$0.0016 0.2181$\pm$0.0009
+Multimodal x x 0.7558$\pm$0.0007 0.4073$\pm$0.0027 0.8362$\pm$0.0006 0.4570$\pm$0.0032
+[…]
+Ours (ResNet) x x 0.7546$\pm$0.0039 0.4438$\pm$0.0053 0.8424$\pm$0.0012 0.4671$\pm$0.0014
+Ours (FT-Trans) x x 0.7610$\pm$0.0015 0.4507$\pm$0.0007 0.8486$\pm$0.0011 0.4796$\pm$0.0027
+Clinical BERT as text encoder
+Text encoder only x 0.6614$\pm$0.0007 0.3080$\pm$0.0009 0.7240$\pm$0.0001 0.2928$\pm$0.0008
+Multimodal x x 0.7584$\pm$0.0005 0.4218$\pm$0.0019 0.8404$\pm$0.0004 0.4686$\pm$0.0026
+[…]
+Ours (ResNet) x x 0.7580$\pm$0.0023 0.4472$\pm$0.0042 0.8474$\pm$0.0006 0.4899$\pm$0.0015
+Ours (FT-Trans) x x 0.7672$\pm$0.0032 0.4541$\pm$0.0032 0.8534$\pm$0.0012 0.4977$\pm$0.0011
+Table 2: Comparison of predictive accuracy on mortality prediction. The best results among models using the same text encoder are in bold, and the overall best results are shaded in grey.
+[…]
+Random Forest x 0.6680$\pm$0.0025 0.3492$\pm$0.0038 0.7270$\pm$0.0018 0.2980$\pm$0.0029
+MLP x 0.6845$\pm$0.0009 0.3736$\pm$0.0011 0.7528$\pm$0.0010 0.3353$\pm$0.0010
+ResNet x 0.6851$\pm$0.0017 0.3774$\pm$0.0018 0.7532$\pm$0.0004 0.3366$\pm$0.0021
+FT-Transformer x 0.6956$\pm$0.0011 0.3790$\pm$0.0021 0.7674$\pm$0.0006 0.3404$\pm$0.0021
+BERT as text encoder
+Text encoder only x 0.6087$\pm$0.0011 0.2960$\pm$0.0016 0.6507$\pm$0.0015 0.2262$\pm$0.0009
+Multimodal x x 0.6903$\pm$0.0018 0.3739$\pm$0.0012 0.7625$\pm$0.0006 0.3427$\pm$0.0007
+[…]
+Ours (ResNet) x x 0.6931$\pm$0.0013 0.3887$\pm$0.0040 0.7655$\pm$0.0011 0.3584$\pm$0.0007
+Ours (FT-Trans) x x 0.7005$\pm$0.0010 0.3809$\pm$0.0024 0.7725$\pm$0.0006 0.3524$\pm$0.0011
+BioBERT as text encoder
+Text encoder only x 0.6057$\pm$0.0016 0.2940$\pm$0.0012 0.6487$\pm$0.0014 0.2214$\pm$0.0012
+Multimodal x x 0.6905$\pm$0.0022 0.3731$\pm$0.0025 0.7606$\pm$0.0007 0.3364$\pm$0.0009
+[…]
+Ours (ResNet) x x 0.6896$\pm$0.0011 0.3918$\pm$0.0037 0.7647$\pm$0.0009 0.3532$\pm$0.0010
+Ours (FT-Trans) x x 0.6986$\pm$0.0011 0.3847$\pm$0.0030 0.7701$\pm$0.0012 0.3504$\pm$0.0014
+Clinical BERT as text encoder
+Text encoder only x 0.6420$\pm$0.0016 0.3261$\pm$0.0015 0.6946$\pm$0.0009 0.2661$\pm$0.0006
+Multimodal x x 0.6933$\pm$0.0014 0.3714$\pm$0.0023 0.7648$\pm$0.0009 0.3544$\pm$0.0018
+[…]
+Ours (ResNet) x x 0.6974$\pm$0.0011 0.4019$\pm$0.0015 0.7705$\pm$0.0009 0.3639$\pm$0.0010
+Ours (FT-Trans) x x 0.7027$\pm$0.0011 0.3942$\pm$0.0020 0.7743$\pm$0.0006 0.3575$\pm$0.0014
+Table 3: Comparison of predictive accuracy on PLOS prediction. The best results among models using the same text encoder are in bold, and the overall best results are shaded in grey.
+[…]
+Random Forest x 0.1891$\pm$0.0006 0.5639$\pm$0.0015
+MLP x 0.1677$\pm$0.0015 0.4937$\pm$0.0037
+ResNet x 0.1654$\pm$0.0006 0.4895$\pm$0.0015
+FT-Transformer x 0.1648$\pm$0.0014 0.4832$\pm$0.0037
+BERT as text encoder
+Text encoder only x 0.2302$\pm$0.0070 0.6523$\pm$0.0159
+Multimodal x x 0.1691$\pm$0.0030 0.4992$\pm$0.0077
+[…]
+Ours (ResNet) x x 0.1435$\pm$0.0052 0.4296$\pm$0.0133
+Ours (FT-Trans) x x 0.1450$\pm$0.0049 0.4345$\pm$0.0129
+BioBERT as text encoder
+Text encoder only x 0.2214$\pm$0.0021 0.6326$\pm$0.0048
+Multimodal x x 0.1725$\pm$0.0028 0.5084$\pm$0.0088
+[…]
+Ours (ResNet) x x 0.1392$\pm$0.0061 0.4187$\pm$0.0157
+Ours (FT-Trans) x x 0.1358$\pm$0.0014 0.4095$\pm$0.0045
+Clinical BERT as text encoder
+Text encoder only x 0.2132$\pm$0.0030 0.6086$\pm$0.0070
+Multimodal x x 0.1606$\pm$0.0022 0.4792$\pm$0.0046
+[…]
+Ours (ResNet) x x 0.1387$\pm$0.0043 0.4178$\pm$0.0110
+Ours (FT-Trans) x x 0.1390$\pm$0.0043 0.4192$\pm$0.0112
+Table 4: Comparison of reliability performance on mortality prediction.
+[…]
+Random Forest x 0.2350$\pm$0.0009 0.6649$\pm$0.0018
+MLP x 0.1950$\pm$0.0018 0.5739$\pm$0.0044
+ResNet x 0.1929$\pm$0.0012 0.5702$\pm$0.0027
+FT-Transformer x 0.1947$\pm$0.0031 0.5704$\pm$0.0079
+BERT as text encoder
+Text encoder only x 0.2316$\pm$0.0043 0.6557$\pm$0.0091
+Multimodal x x 0.2017$\pm$0.0028 0.5960$\pm$0.0082
+[…]
+Ours (ResNet) x x 0.1815$\pm$0.0045 0.5398$\pm$0.0107
+Ours (FT-Trans) x x 0.1927$\pm$0.0039 0.5617$\pm$0.0100
+BioBERT as text encoder
+Text encoder only x 0.2297$\pm$0.0029 0.6518$\pm$0.0063
+Multimodal x x 0.1974$\pm$0.0026 0.5888$\pm$0.0043
+[…]
+Ours (ResNet) x x 0.1752$\pm$0.0048 0.5241$\pm$0.0123
+Ours (FT-Trans) x x 0.1832$\pm$0.0045 0.5376$\pm$0.0113
+Clinical BERT as text encoder
+Text encoder only x 0.2225$\pm$0.0031 0.6350$\pm$0.0069
+Multimodal x x 0.2027$\pm$0.0056 0.5871$\pm$0.0136
+[…]
+Ours (ResNet) x x 0.1694$\pm$0.0018 0.5082$\pm$0.0042
+Ours (FT-Trans) x x 0.1760$\pm$0.0018 0.5205$\pm$0.0038
+Table 5: Comparison of reliability performance on PLOS prediction.
 Overall, our framework demonstrated superior performance across both tasks.
-In the mortality prediction task, our framework using MLP and Clinical BERT as the backbone achieved the highest F1 score (0.4629) and the lowest Brier score (0.1176) and NLL (0.3594).
-Furthermore, with FT-Transformer and Clinical BERT, it achieved the highest balanced accuracy (BACC) of 0.7672, AUROC of 0.8534, and AUPRC of 0.4977.
+In the mortality prediction task, our framework using MLP and Clinical BERT as the backbones achieved the highest F1 score (0.4629).
+Furthermore, with FT-Transformer and Clinical BERT, it achieved the highest BACC of 0.7672, AUROC of 0.8534, and AUPRC of 0.4977.
 Compared to the best baseline models, our framework improved predictive performance by approximately 1.05% in BACC, 9.74% in F1 score, 1.28% in AUROC, and 6.21% in AUPRC.
-Additionally, it significantly enhanced prediction reliability, with reductions of about 26.8% in Brier score and 25.0% in NLL.
-In the PLOS prediction task, our framework demonstrated exceptional performance across all key predictive evaluation metrics, achieving a BACC of 0.7027, an F1 score of 0.4019, an AUROC of 0.7743, and an AUPRC of 0.3639.
-Compared to the best baseline model, our framework showed notable improvements: a 1.02% increase in BACC, a 6.04% increase in F1 score, a 0.9% increase in AUROC, and a 2.68% increase in AUPRC.
-Additionally, it enhanced prediction reliability with a 15.1% reduction in Brier score and a 13.3% reduction in NLL.
-Models that relied solely on structured data or free-text notes underperformed compared to our proposed framework in both predictive and reliability evaluation metrics.
-While the multimodal approach showed marginally better performance than other baselines in the mortality prediction task, it delivered weaker results in the PLOS prediction task.
-Evaluation on different fusion settings
-Figure 3 illustrates the evaluation of our framework across three fusion settings applied to different backbone models for mortality prediction.
-For the corresponding results in the PLOS prediction task, please refer to Appendix F.
-Both tasks exhibited similar patterns in the analysis.
-Regarding predictive evaluation metrics, the models achieved comparable BACC across the three fusion settings for both tasks.
+In the PLOS prediction task, our framework demonstrated exceptional performance across all evaluation metrics for predictive accuracy , achieving a BACC of 0.7027, an F1 score of 0.4019, an AUROC of 0.7743, and an AUPRC of 0.3639.
+Compared to the best baselines, our framework showed notable improvements: a 1.02% increase in BACC, a 6.04% increase in F1 score, a 0.9% increase in AUROC, and a 2.68% increase in AUPRC.
+Over the baseline models, the multimodal approaches showed marginally worse predictive accuracy than others in both tasks.
+Specifically, in the mortality prediction task, the multimodal approach with clinical BERT as the backbone achieved the highest F1 of 0.4218 and AUPRC of 0.4686.
+In the PLOS task, it achieved the highest AUPRC of 0.3544.
+5.2.2 Reliability
+In clinical settings, assessing prediction reliability is as important as evaluating predictability.
+Table 4 and 5 present the comparison of model performance based on reliability across both prediction tasks.
+In the mortality prediction task, our framework achieved the lowest Brier score (0.1176) and NLL (0.3594) with MLP and Clinical BERT as the backbones.
+It improved prediction reliability significantly compared to the best baseline models, reducing the Brier score by about 26.8% and NLL by 25.0%.
+In the PLOS prediction task, our framework also showed strong performance, with a Brier score of 0.1637 and an NLL of 0.4943.
+These results showed notable improvements over the best baseline model, with a 15.1% reduction in Brier score and a 13.3% reduction in NLL.
+Compared to other baselines, the multimodal approaches exhibited slightly better reliability in the mortality prediction task, achieving a Brier score of 0.1606 and an NLL of 0.4792.
+However, in the PLOS prediction tasks, the multimodal approaches demonstrated weaker reliability.
+Figure 3: The illustration of different fusion settings.
+5.3 Evaluation on different fusion settings
+We also validate the effectiveness of our fusion setting in EHR groups by comparing it with two other additional fusion settings: data types and data sources.
+The illustration is shown in Figure 3.
+For the data types fusion, we divided the structured data into two types: numerical and categorical data.
+For the data sources fusion, we split the structured data into four sources: demographics, vital signs/lab tests, medical treatments, and comorbidities, and categorized the clinical notes into four types: Nursing, Nursing/Other, Physician, and Radiology notes.
+Figures 4 and 5 illustrate the evaluation of our framework using clinical BERT as text encoder across three fusion settings for both tasks.
+Evaluation results about models using BERT and BioBERT as text encoders can be found in Appendix A.
+Regarding predictive accuracy, the models achieved lower BACC across the three fusion settings for both tasks.
 Notably, the model based on modalities outperformed those based on data types and data sources in F1, AUROC, and AUPRC metrics.
-In terms of reliability evaluation metrics, the model based on modalities demonstrated superior reliability across different backbone models.
-Figure 3: The evaluation of our framework on different fusion settings for mortality prediction: (1) modalities, (2) data types, (3) data sources.
-DISCUSSION
-Principle findings
-In this section, we present our findings across four key areas: the advantages of using free-text notes for mortality and PLOS prediction, the effectiveness of evidence-based multimodal framework, the impact of various fusion settings, and the influence of encoder selection.
-Benefits of free-text notes
-From Table 2, it is evident that integrating free-text notes with structured data enhances ICU outcome prediction, boosting predictive performance and reliability.
+In terms of reliability evaluation metrics, the model based on modalities demonstrated superior reliability across three different structured data encoders.
+Additionally, Figure 4 and 5 suggest a positive relationship between predictability and reliability in our framework.
+Figure 4: The evaluation of our framework using Clinical BERT as text encoder on different fusion settings for mortality prediction: (1) modalities, (2) data types, (3) data sources.
+Figure 5: The evaluation of our framework using Clinical BERT as text encoder on different fusion settings for PLOS prediction: (1) modalities, (2) data types, (3) data sources.
+6 DISCUSSION
+In this section, we present our findings across four key areas: the advantages of using free-text notes for mortality and PLOS prediction, the effectiveness of the evidence-based multimodal framework, the impact of various fusion settings, and the influence of encoder selection.
+6.1 Benefits of free-text notes
+From Tables 2 and 3, it is evident that integrating free-text notes with structured data enhances ICU outcome prediction, boosting predictability.
 Free-text notes provide information not included in structured EHR data, such as nursing details, physician documentation, and radiology reports after ICU admission.
 This suggests that free-text EHR notes and structured inputs can complement each other in predictive modeling, leading to improved performance.
-Effectiveness of evidence-based multimodal framework
-Our evidence-based framework demonstrated improved predictive performance and higher reliability compared to existing multimodal approaches.
-This improvement is likely due to the heterogeneous nature of EHRs, which often contain imprecise information from diverse sources.
-By effectively extracting and integrating informative evidence through modalities, our framework addresses these challenges.
-Additionally, the current multimodal approach achieved higher recall in both tasks (refer to Appendix E).
-High recall reflects the model’s ability to capture the most true positive instances, often at the cost of accepting more false positives.
-In ICU settings, this could result in the unnecessary utilization of medical resources and equipment.
-In contrast, our framework strikes a better balance by demonstrating higher precision and specificity, reducing false positives.
-This capability is critical to appropriately allocate medical resources to ICU patients in real need.
-It is also noteworthy that the improvement in BACC (1.05% for mortality prediction and 1.02% for PLOS prediction) is smaller than the improvement in F1 score (9.74% for mortality prediction and 6.04% for PLOS prediction).
-The F1 score, which emphasizes the balance between precision and recall, is particularly valuable for assessing improvements in identifying positive samples within imbalanced datasets, especially when positive outcomes are rare.
-Given that ICU outcomes are a minority in our dataset, this suggests that our framework is particularly effective at correctly identifying these critical outcomes.
-Impacts of different fusion settings
-To explore the performance of our framework across different fusion settings, it is evident that models based on modalities achieved higher F1 scores but lower BACCs compared to the other two fusion settings.
-This discrepancy arises from the metrics’ focus: the F1 score emphasizes performance on the positive class by balancing precision and recall, while balanced accuracy (BACC) provides an overall assessment of recall across all classes.
-Thus, models based on modalities are more effective at identifying ICU outcomes, likely due to the enhanced integration of information from independent sources facilitated by evidence theory.
-This observation aligns with our analysis of data independence.
-Figures 4 and 5 visualize the independence of structured features and four types of free-text notes, respectively.
-In Figure 4, correlation coefficients confirm that features within structured data are not independent.
-Meanwhile, Figure 5 shows that Radiology notes form a distinct cluster, while the other types of notes overlap, indicating a lack of independence among them.
-Additionally, our analysis highlights a general relationship between predictive performance and reliability.
-Figure 4: The pair plots and correlation coefficients of some structured features from different sources.
-Figure 5: The t-SNE visualization on extracted embeddings of four different types of free-texts in EHRs: Nursing, Nursing/other, Physician, Radiology.
-Influence of encoders selection
-The evaluations presented in Tables 2 and 3 reveal that the choice of encoders plays a critical role in determining predictive performance.
-Among the models assessed, the FT-Transformer stands out as highly effective for tabular data, aligning with the findings of 23.
-This effectiveness can be attributed to the transformer’s capability to capture complex relationships among transformed numerical and categorical features, which enhances its predictive power.
-For pre-trained language encoders, Clinical BERT demonstrates the best performance in extracting clinical information from free-text notes.
-Its superiority stems from its unique pre-training on clinical text from the MIMIC-III database.
-This specialized pre-training enables Clinical BERT to generate more informative embeddings by leveraging its pre-learned clinical knowledge and domain-specific term embeddings, resulting in improved model performance.
-CONCLUSION
-In this paper, we address the challenge of predicting ICU outcomes by introducing an evidence-based multimodal framework that models both structured EHR data and free-text EHR notes.
-Our framework transforms deep features extracted from these two modalities into evidence through evidence mapping module, which is then fused using Dempster’s rule to make final predictions.
-Through experiments on the MIMIC-III dataset, we demonstrate the effectiveness of our framework in managing heterogeneous multimodal EHR data, reducing false positives and potentially improving the allocation of medical resources in the ICU.
-While this paper focuses on binary classification tasks, many clinical applications require solutions for multiclass tasks (e.g., disease diagnosis) and continuous regression tasks (e.g., survival prediction).
-These are equally important and relevant for advancing clinical practice.
-In the future, we plan to expand our framework by incorporating additional data modalities, such as time series and medical images, to provide deeper clinical insights and enhance model performance.
-We also aim to extend the framework to handle multimodal EHR multiclass tasks, offering valuable predictive guidance for complex clinical scenarios.
-Additionally, we intend to investigate regression tasks, leveraging the recently introduced Epistemic Random Fuzzy Set (ERFS) theory 32, 33 and further building on developments in evidential regression 34.
-ACKNOWLEDGMENTS
-This research is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG-GC-2019-001-2B).
-APPENDIX
-A  Dempster-Shafer (DS) Theory
-Let $\Omega=\{\omega_{1},\omega_{2},\cdots,\omega_{M}\}$ be a finite set of hypotheses about some question, called the frame of discernment.
-Evidence about a variable taking values in $\Omega$ can be represented by a mass function: $2^{\Omega}$ to [0,1] such that $m(\emptyset)=0$ and
-$\sum_{A\subseteq\Omega}m(A)=1.$ (1)
-For any hypothesis $A\subseteq\Omega$, the quantity $m(A)$ is interpreted as a share of a unit mass of belief allocated to the hypothesis that the truth is in $A$, and which cannot be allocated to any strict subset of $A$ based on the available evidence.
-Set $A$ is called a focal set of $m$ if $m(A)>0$.
-A mass function is said to be Bayesian if its focal sets are singletons, and logical if it has only one focal set.
-Two mass functions $m_{1}$ and $m_{2}$ representing independent items of evidence can be combined conjunctively by Dempster’s combination rule 17 $\oplus$ as
-$(m_{1}\oplus m_{2})(A)=\frac{\sum_{B\cap C=A}m_{1}(B)m_{2}(C)}{1-\sum_{B\cap C=\emptyset}m_{1}(B)m_{2}(C)},$ (2)
-for all $A\neq\emptyset$, where $\sum_{B\cap C=\emptyset}m_{1}(B)m_{2}(C)$ is the degree of conflict among the two pieces of evidence, The nice information fusion attribute of DS theory points out the high potential of using DS theory for heterogenetic medical data analysis.
-After aggregating all available evidence, we adopt the pignistic transformation, as proposed by Smets in the Transferable Belief Model 35, to make a final decision based on the combined mass function.
-This framework bases decisions on the pignistic probability distribution with the following expression:
-$p(\omega)=\sum_{A\subseteq\Omega:\omega\in A}\frac{m(A)}{|A|},\forall\omega\in\Omega.$ (3)
-B  Evidence mapping (EM)
-As illustrated in Figure 6, the evidence mapping module consists of one input layer, one hidden layer, and one output layer.
-The input layer is composed of $H$ units ($H$ is the number of prototypes), whose weights vectors are prototypes $\pi_{1},\pi_{2},\cdots,\pi_{H}$ in input space.
-The activation of unit $h$ in the input layer is
-$s_{h}=\beta_{h}\mathrm{exp}(-\gamma_{h}d_{h}^{2}),$ (4)
-Figure 6: The illustration of evidence mapping
-where $d_{h}=||x-\pi_{h}||$ denotes the Euclidean distance between input vector $x$ and prototype $\pi_{h}$ , $\gamma_{h}>0$ is a scale parameter, and $\beta_{h}\in[0,1]$ is an extra parameter.
-The hidden layer computes mass functions $m_{h}$ (evidence) of each prototype $\pi_{h}$ is defined as:
-$m_{h}(\{\omega_{c}\})$ $=u_{h}^{(c)}s_{h},\quad c=1,2,\cdots,M,$ (5a)
-$m_{h}(\Omega)$ $=1-s_{h},$ (5b)
-where $u_{h}^{(c)}$ is the membership degree of prototype $h$ to class $\omega_{c}$, $\sum_{c=1}^{M}u_{h}^{(c)}=1$, and $M$ is the number of classes.
-Therefore, the vector of mass functions induced by prototypes is denoted as:
-$m_{h}=(m_{h}(\{\omega_{1}\}),m_{h}(\{\omega_{2}\}),\cdots,m_{h}(\{\omega_{M}\}),m_{h}(\Omega))\in\mathbb{R}^{M+1}.$ (6)
-Finally, the mass functions are then aggregated by Dempster’s combination rule using Eq.
-2 in the output layer.
-A combined mass function $m$ is computed as the orthogonal sum of the $H$ mass functions:
-$m=m_{1}\oplus m_{2}\oplus\cdots\oplus m_{H}\in\mathbb{R}^{M+1}.$ (7)
-The combined mass functions (the outputs of the EM module) represent the degrees of belief about the given class with $m(\{\omega_{c}\})$, as well as its prediction uncertainty with $m(\Omega)$.
-In our binary classification case, the dimension of EM module outputs would be three.
-C  Evidence fusion
-To combine multiple mass functions $m_{1},m_{2},\cdots,m_{K}$ from different modalities/data types/data sources, Dempster’s combination rule is applied again Eq.
-2 to aggregate evidence for multiple sources for final evidence generation.
-$m=m_{1}\oplus m_{2}\oplus\cdots\oplus m_{K}\in\mathbb{R}^{M+1},$ (8)
-where $K$ is the number of mass functions to combine. For example, $K$=2 for the fusion of evidence from structured EHRs and free-text notes.
-D  The details of augmented learning algorithm
-Let $p_{i}=(p_{i}(\omega_{1}),\cdots,p_{i}(\omega_{c}),\cdots,p_{i}(\omega_{M}))$ be the final probability after the pignistic transformation (3) for training sample $i$, and $y_{i}=(y_{i,1},y_{i,2},\cdots,y_{i,M})$ denotes the one-hot encoding for corresponding ground-truth labels.
-The main loss function $\mathcal{L}_{main}$ is computed as:
-$\mathcal{L}_{p}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{c=1}^{M}w_{c}y_{i,c}\mathrm{log}(p_{i}(\omega_{c})),$ (9)
-where $N$ is the number of training samples, $M$ denotes the number of classes, and $w_{c}$ is the weight assigned to each class to address the class imbalance issue.
-Moreover, two auxiliary cross-entropy losses are introduced to optimize the feature representation performance of the encoders, as the evidence mapping module performs more effectively with high-quality representations.
-Firstly, to regulate the representation generated by encoders, we added an additional fully connected network (FCN) to generate logits $o_{i}$ for each modality.
-Let $o^{s}_{i}=(o_{i,1}^{s},o_{i,2}^{s},\cdots,o_{i,M}^{s})$ and $o^{n}_{i}=(o_{i,1}^{n},o_{i,2}^{n},\cdots,o_{i,M}^{n})$ be the logits from the encoders for structured data and free-text notes, respectively, the cross-entropy losses $\mathcal{L}^{s}_{aux}$ and $\mathcal{L}^{n}_{aux}$ are then calculated with $y_{i}$ for structured data and notes, respectively:
-$\mathcal{L}^{s}_{aux}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{c=1}^{M}w_{c}y_{i,c}\mathrm{log}\frac{\mathrm{exp}(o_{i,c}^{s})}{\sum_{b=1}^{M}\mathrm{exp}(o_{i,b}^{s})},$ (10)
-$\mathcal{L}^{n}_{aux}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{c=1}^{M}w_{c}y_{i,c}\mathrm{log}\frac{\mathrm{exp}(o_{i,c}^{n})}{\sum_{b=1}^{M}\mathrm{exp}(o_{i,b}^{n})},$ (11)
-Ultimately, the overall loss function $\mathcal{L}_{overall}$ is defined as follows:
-$\mathcal{L}_{overall}=\mathcal{L}_{main}+\alpha\mathcal{L}^{s}_{aux}+\beta\mathcal{L}^{n}_{aux},$ (12)
-where $\alpha,\beta$ are the hyperparameters that control the balance between the main loss and the auxiliary cross-entropy losses.
-E  The model comparison results in metrics (specificity, recall, sensitivity, and NPV)
+6.2 Effectiveness on multimodal evidence fusion
+Accurate and reliable ICU decision support
+The proposed framework outperforms existing multimodal approaches by leveraging belief function theory for the effective fusion of structured and unstructured EHR data.
+This enables more accurate and robust predictions, which is essential for clinical decision-making in ICU settings.
+Notably, while the improvement in BACC is not significant (1.05% for mortality and 1.02% for PLOS prediction), the F1 score shows a significant increase (9.74% for mortality and 6.04% for PLOS prediction).
+This highlights the ability of the framework to identify critical ICU cases, as the F1 score prioritizes precision and recall, making it especially valuable in imbalanced datasets where positive cases are rare.
+Given the high-stakes nature of ICU outcomes, this improvement suggests our model enhances early identification of high-risk patients, potentially supporting timelier interventions in critical care.
+Unlike traditional multimodal approaches, our evidence-based framework demonstrates greater prediction reliability, which is particularly valuable in clinical decision-making.
+By effectively handling uncertainty and inconsistencies in patient data under the proposed multimodal fusion framework, our approach ensures more trustworthy predictions with lower Brier and NLL (shown in Table 4 and 5).
+This enhanced reliability is crucial for ICU decision support, where inaccurate predictions can lead to overuse of critical resources or missed early interventions.
+Efficient resource allocation
+The experimental results in Tables 6 and 7 demonstrate that while existing multimodal approaches can effectively capture true positive instances, they often come at the cost of increased false positives.
+In ICU settings, such false positives can lead to the unnecessary use of medical resources and equipment.
+In contrast, our framework achieves a better balance by demonstrating higher precision and specificity, effectively reducing false positives.
+This capability is crucial for ensuring that ICU resources are allocated appropriately to patients in critical need.
 […]
@@
 FT-Transformer x 0.2838$\pm$0.0015 0.7832$\pm$0.0054 0.7352$\pm$0.0019 0.9620$\pm$0.0009
-BERT x 0.1767$\pm$0.0030 0.6966$\pm$0.0225 0.5634$\pm$0.0217 0.9330$\pm$0.0021
-BioBERT x 0.1793$\pm$0.0011 0.6029$\pm$0.0278 0.6299$\pm$0.0194 0.9225$\pm$0.0028
-Clinical Bert x 0.1962$\pm$0.0012 0.7158$\pm$0.0080 0.6071$\pm$0.0073 0.9410$\pm$0.0009
-Multimodal (BERT) x x 0.2761$\pm$0.0034 0.7807$\pm$0.0063 0.7255$\pm$0.0064 0.9611$\pm$0.0008
-Multimodal (BioBERT) x x 0.2741$\pm$0.0031 0.7934$\pm$0.0065 0.7182$\pm$0.0067 0.9629$\pm$0.0008
-Multimodal (Clinical Bert) x x 0.2908$\pm$0.0027 0.7678$\pm$0.0063 0.7489$\pm$0.0054 0.9601$\pm$0.0008
-Ours (MLP-BERT) x x 0.3297$\pm$0.0076 0.6957$\pm$0.0168 0.8094$\pm$0.0114 0.9522$\pm$0.0020
-Ours (MLP-BioBERT) x x 0.3377$\pm$0.0115 0.6792$\pm$0.0218 0.8191$\pm$0.0148 0.9503$\pm$0.0023
-Ours (MLP-Clinical Bert) x x 0.3609$\pm$0.0075 0.6478$\pm$0.0123 0.8454$\pm$0.0079 0.9472$\pm$0.0013
-Ours (ResNet-BERT) x x 0.3179$\pm$0.0073 0.7206$\pm$0.0168 0.7916$\pm$0.0118 0.9550$\pm$0.0020
-Ours (ResNet-BioBERT) x x 0.3244$\pm$0.0100 0.7092$\pm$0.0219 0.8000$\pm$0.0150 0.9538$\pm$0.0025
-Ours (ResNet-Clinical Bert) x x 0.326$\pm$0.0067 0.7151$\pm$0.0131 0.8010$\pm$0.0097 0.9546$\pm$0.0015
-Ours (FT-Trans-BERT) x x 0.3163$\pm$0.0080 0.7433$\pm$0.0109 0.7836$\pm$0.0110 0.9580$\pm$0.0012
-Ours (FT-Trans-BioBERT) x x 0.3282$\pm$0.0015 0.7193$\pm$0.0055 0.8027$\pm$0.0027 0.9553$\pm$0.0007
-Ours (FT-Trans-Clinical Bert) x x 0.3289$\pm$0.0070 0.7371$\pm$0.0171 0.7974$\pm$0.0108 0.9578$\pm$0.0021
-(a) Mortality prediction
+BERT as text encoder
+Text encoder only x 0.1767$\pm$0.0030 0.6966$\pm$0.0225 0.5634$\pm$0.0217 0.9330$\pm$0.0021
+Multimodal x x 0.2761$\pm$0.0034 0.7807$\pm$0.0063 0.7255$\pm$0.0064 0.9611$\pm$0.0008
+[…]
+Ours (ResNet) x x 0.3179$\pm$0.0073 0.7206$\pm$0.0168 0.7916$\pm$0.0118 0.9550$\pm$0.0020
+Ours (FT-Trans) x x 0.3163$\pm$0.0080 0.7433$\pm$0.0109 0.7836$\pm$0.0110 0.9580$\pm$0.0012
+BioBERT as text encoder
+Text encoder only x 0.1793$\pm$0.0011 0.6029$\pm$0.0278 0.6299$\pm$0.0194 0.9225$\pm$0.0028
+Multimodal x x 0.2741$\pm$0.0031 0.7934$\pm$0.0065 0.7182$\pm$0.0067 0.9629$\pm$0.0008
+[…]
+Ours (ResNet) x x 0.3244$\pm$0.0100 0.7092$\pm$0.0219 0.8000$\pm$0.0150 0.9538$\pm$0.0025
+Ours (FT-Trans) x x 0.3282$\pm$0.0015 0.7193$\pm$0.0055 0.8027$\pm$0.0027 0.9553$\pm$0.0007
+Clinical BERT as text encoder
+Text encoder only x 0.1962$\pm$0.0012 0.7158$\pm$0.0080 0.6071$\pm$0.0073 0.9410$\pm$0.0009
+Multimodal x x 0.2908$\pm$0.0027 0.7678$\pm$0.0063 0.7489$\pm$0.0054 0.9601$\pm$0.0008
+[…]
+Ours (ResNet) x x 0.3260$\pm$0.0067 0.7151$\pm$0.0131 0.8010$\pm$0.0097 0.9546$\pm$0.0015
+Ours (FT-Trans) x x 0.3289$\pm$0.0070 0.7371$\pm$0.0171 0.7974$\pm$0.0108 0.9578$\pm$0.0021
+Table 6: Comparison of class-specific prediction accuracy on mortality prediction.
 […]
@@
 FT-Transformer x 0.2603$\pm$0.0026 0.6974$\pm$0.0061 0.6938$\pm$0.0067 0.9369$\pm$0.0007
-BERT x 0.2004$\pm$0.0031 0.5685$\pm$0.0147 0.6488$\pm$0.0160 0.9070$\pm$0.0009
-BioBERT x 0.2006$\pm$0.0022 0.5504$\pm$0.0160 0.6610$\pm$0.0141 0.9051$\pm$0.0013
-Clinical Bert x 0.2208$\pm$0.0024 0.6251$\pm$0.0133 0.6587$\pm$0.0116 0.9193$\pm$0.0014
-Multimodal (BERT) x x 0.2568$\pm$0.0029 0.6892$\pm$0.0141 0.6915$\pm$0.0108 0.9347$\pm$0.0022
-Multimodal (BioBERT) x x 0.2554$\pm$0.0034 0.6943$\pm$0.0136 0.6868$\pm$0.0114 0.9358$\pm$0.0018
-Multimodal (Clinical Bert) x x 0.2507$\pm$0.0045 0.7201$\pm$0.0188 0.6664$\pm$0.0163 0.9394$\pm$0.0024
-Ours (MLP-BERT) x x 0.2809$\pm$0.0038 0.6218$\pm$0.0050 0.7538$\pm$0.0064 0.9281$\pm$0.0005
-Ours (MLP-BioBERT) x x 0.2892$\pm$0.0048 0.6022$\pm$0.0061 0.7710$\pm$0.0071 0.9262$\pm$0.0007
-Ours (MLP-Clinical Bert) x x 0.2980$\pm$0.0075 0.5899$\pm$0.0134 0.7841$\pm$0.0117 0.9254$\pm$0.0012
-Ours (ResNet-BERT) x x 0.2787$\pm$0.0071 0.6460$\pm$0.0158 0.7402$\pm$0.0148 0.9314$\pm$0.0016
-Ours (ResNet-BioBERT) x x 0.2848$\pm$0.0059 0.6212$\pm$0.0126 0.7580$\pm$0.0119 0.9285$\pm$0.0012
-Ours (ResNet-Clinical Bert) x x 0.2968$\pm$0.0029 0.6227$\pm$0.0073 0.7720$\pm$0.0058 0.9299$\pm$0.0008
-Ours (FT-Trans-BERT) x x 0.2595$\pm$0.0036 0.7180$\pm$0.0116 0.6830$\pm$0.0110 0.9402$\pm$0.0014
-Ours (FT-Trans-BioBERT) x x 0.2668$\pm$0.0042 0.6914$\pm$0.0106 0.7059$\pm$0.0109 0.9368$\pm$0.0012
-Ours (FT-Trans-Clinical Bert) x x 0.2782$\pm$0.0026 0.6766$\pm$0.0046 0.7288$\pm$0.0050 0.9359$\pm$0.0005
-(b) PLOS prediction
-Table 3: The model comparison for mortality (a) and PLOS prediction (b) tasks in metrics (precision, recall, specificity, and NPV).
-F  The evaluation on different fusion settings for PLOS prediction
-Figure 7: The evaluation of our framework on different fusion settings for PLOS prediction: (1) modalities, (2) data types, (3) data sources.
+BERT as text encoder
+Text encoder only x 0.2004$\pm$0.0031 0.5685$\pm$0.0147 0.6488$\pm$0.0160 0.9070$\pm$0.0009
+Multimodal x x 0.2568$\pm$0.0029 0.6892$\pm$0.0141 0.6915$\pm$0.0108 0.9347$\pm$0.0022
+[…]
+Ours (ResNet) x x 0.2787$\pm$0.0071 0.6460$\pm$0.0158 0.7402$\pm$0.0148 0.9314$\pm$0.0016
+Ours (FT-Trans) x x 0.2595$\pm$0.0036 0.7180$\pm$0.0116 0.6830$\pm$0.0110 0.9402$\pm$0.0014
+BioBERT as text encoder
+Text encoder only x 0.2006$\pm$0.0022 0.5504$\pm$0.0160 0.6610$\pm$0.0141 0.9051$\pm$0.0013
+Multimodal x x 0.2554$\pm$0.0034 0.6943$\pm$0.0136 0.6868$\pm$0.0114 0.9358$\pm$0.0018
+[…]
+Ours (ResNet) x x 0.2848$\pm$0.0059 0.6212$\pm$0.0126 0.7580$\pm$0.0119 0.9285$\pm$0.0012
+Ours (FT-Trans) x x 0.2668$\pm$0.0042 0.6914$\pm$0.0106 0.7059$\pm$0.0109 0.9368$\pm$0.0012
+Clinical BERT as text encoder
+Text encoder only x 0.2208$\pm$0.0024 0.6251$\pm$0.0133 0.6587$\pm$0.0116 0.9193$\pm$0.0014
+Multimodal x x 0.2507$\pm$0.0045 0.7201$\pm$0.0188 0.6664$\pm$0.0163 0.9394$\pm$0.0024
+[…]
+Ours (ResNet) x x 0.2968$\pm$0.0029 0.6227$\pm$0.0073 0.7720$\pm$0.0058 0.9299$\pm$0.0008
+Ours (FT-Trans) x x 0.2782$\pm$0.0026 0.6766$\pm$0.0046 0.7288$\pm$0.0050 0.9359$\pm$0.0005
+Table 7: Comparison of class-specific prediction accuracy on PLOS prediction.
+6.3 Analysis on different fusion settings
+To explore the performance of our framework across different fusion settings, it is evident that models based on modalities achieved higher F1 scores but lower BACCs compared to the other two fusion settings.
+This discrepancy arises from the metrics’ focus: the F1 score emphasizes performance on the positive class by balancing precision and recall, while BACC provides an overall assessment of recall across all classes.
+Thus, models based on modalities are more effective at identifying ICU outcomes, likely due to the enhanced integration of information from independent sources facilitated by belief function theory.
+This observation aligns with our analysis of data independence.
+Figures 6 and 7 visualize the independence of structured features and four types of free-text notes, respectively.
+In Figure 6, correlation coefficients confirm that features within structured data are not independent.
+Meanwhile, Figure 7 shows that Radiology notes form a distinct cluster, while the other types of notes overlap, indicating a lack of independence among them.
+Moreover, the difference in BACC performance using the FT-Transformer across three fusion settings is smaller than with the MLP.
+This is likely due to the stronger predictive capability of the FT-Transformer.
+Figure 6: The pair plots and correlation coefficients of some structured features from different sources.
+Figure 7: The t-SNE visualization on extracted embeddings of four different types of free-texts in EHRs: Nursing, Nursing/other, Physician, Radiology.
+6.4 Influence of encoders selection
+The evaluations presented in Table 2 and 3 reveal that the choice of encoders plays a critical role in determining predictability.
+Among the assessed models, the FT-Transformer stands out as highly effective for structured tabular data, aligning with the findings of [34].
+This effectiveness can be attributed to the transformer’s capability to capture complex relationships among transformed numerical and categorical features, which enhances its predictive power.
+For pre-trained language encoders, Clinical BERT demonstrates the best performance in extracting clinical information from free-text notes.
+Its superiority stems from its unique pre-training on clinical text from the MIMIC-III database.
+This specialized pre-training enables Clinical BERT to generate more informative embeddings by leveraging its pre-learned clinical knowledge and domain-specific term embeddings, resulting in improved model performance.
+7 CONCLUSION
+In this paper, we address the challenge of accurately and reliably predicting ICU outcomes by introducing a multimodal framework based on belief function theory that models both structured EHR data and free-text EHR notes.
+Our framework transforms deep features extracted from these two modalities into evidence through the evidence mapping module, which is then fused using Dempster’s rule to make final predictions.
+Through experiments on the MIMIC-III dataset, we demonstrate the effectiveness of our framework in terms of predictive accuracy and reliability.
+The study highlights its capability in managing heterogeneous multimodal EHR data, reducing false positives and potentially improving the allocation of medical resources in the ICU.
+While this paper focuses on binary classification tasks, many clinical applications require solutions for multiclass tasks (e.g., disease diagnosis) and continuous regression tasks (e.g., survival prediction).
+These are equally important and relevant for advancing clinical practice.
+In the future, we plan to expand our framework by incorporating additional data modalities, such as time series and medical images, to provide deeper clinical insights and enhance model performance.
+We also aim to extend the framework to handle multimodal EHR multiclass tasks, offering valuable predictive guidance for complex clinical scenarios.
+Additionally, we intend to investigate regression tasks, leveraging the recently introduced Epistemic Random Fuzzy Set (ERFS) theory [46, 26] and further building on developments in evidential regression [24].
+[…]
+Acknowledgements
+This research is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG-GC-2019-001-2B).
+[…]
+Code Availability The code implementation is available on Github repository (https://github.com/yuchengruan/evid_multimodal_ehr).
+Declarations
+[…]
+[…]
+Appendix A Evaluation on different fusion settings using BERT and BioBERT as text encoders
+Figure 8: The evaluation of our framework using BERT as text encoder on different fusion settings for mortality prediction: (1) modalities, (2) data types, (3) data sources.
+Figure 9: The evaluation of our framework using BERT as text encoder on different fusion settings for PLOS prediction: (1) modalities, (2) data types, (3) data sources.
+Figure 10: The evaluation of our framework using BioBERT as text encoder on different fusion settings for mortality prediction: (1) modalities, (2) data types, (3) data sources.
+Figure 11: The evaluation of our framework using BioBERT as text encoder on different fusion settings for PLOS prediction: (1) modalities, (2) data types, (3) data sources.